### v1

# Multi-Domain Evolutionary Optimization of Network Structures

Abstract
Multi-Task Evolutionary Optimization (MTEO), an important field focusing on addressing complex problems through optimizing multiple tasks simultaneously, has attracted much attention. While MTEO has been primarily focusing on task similarity, there remains a hugely untapped potential in harnessing the shared characteristics between different domains to enhance evolutionary optimization. For example, real-world complex systems usually share the same characteristics, such as the power-law rule, small-world property and community structure, thus making it possible to transfer solutions optimized in one system to another to facilitate the optimization. Drawing inspiration from this observation of shared characteristics within complex systems, we set out to extend MTEO to a novel framework—multi-domain evolutionary optimization (MDEO). To examine the performance of the proposed MDEO, we utilize a challenging combinatorial problem of great security concern—community deception in complex networks as the optimization task. To achieve MDEO, we propose a community-based measurement of graph similarity to manage the knowledge transfer among domains. Furthermore, we develop a graph representation-based network alignment model that serves as the conduit for effectively transferring solutions between different domains. Moreover, we devise a self-adaptive mechanism to determine the number of transferred solutions from different domains and introduce a novel mutation operator based on the learned mapping to facilitate the utilization of knowledge from other domains. Experiments on eight real-world networks of different domains demonstrate MDEO’s superiority in efficacy compared to classical evolutionary optimization. Simulations of attacks on the community validate the effectiveness of the proposed MDEO in safeguarding community security.
Index Terms:
Complex network, multi-domain evolutionary optimization, multi-task evolutionary optimization, transfer learning.
I Introduction
Multi-task evolutionary optimization (MTEO) has stood as an emerging framework to solve complex optimization problems, undergoing remarkable advancements [1, 2, 3, 4] since the multi-factorial evolutionary algorithm (MFEA) was proposed [5]. MTEO aims to address situations where there are multiple related tasks or objectives being optimized simultaneously, with the goal of achieving better performance on each task.
While most existing endeavors primarily concentrate on exploring relationships between tasks, a research gap persists regarding correlations between different domains in evolutionary optimization. In real-world applications, we encounter a multitude of networks that capture different aspects of complex systems [6]. These networks may represent social networks [7] or transportation networks [8]. Many of these networks share common structural properties, such as power-law distribution [9, 10], community structure [11, 12] and small-world characteristic [13, 14, 15].
In real-world scenarios, different domains often have common objective, such as the task of critical node mining in complex networks [16, 17], which can contribute to the prevention of catastrophic outages in the power networks [18], the identification of drug target candidates in protein networks [19], or the improvement of the robustness in communication networks [20]. The existence of shared attributes and structural properties across real-world networks presents an avenue for knowledge transfer and collaborative learning when optimizing an identical task. To date, multi-task evolutionary optimization of complex networks has already been extensively studied [21, 22], however, leveraging the correlation of domains to facilitate the optimization of the network structure remains underexplored. Inspired by MTEO, we propose a new framework called multi-domain evolutionary optimization (MDEO) for network structures, in which solutions obtained from optimizing one network can be effectively transferred and adapted to optimize another network with improved efficacy. The following are the main differences between MTEO and MDEO, justifying the novelty of our work:
Granularity of Tasks: In MTEO, the emphasis is on optimizing different tasks, which may be distinct and well-defined within a single network. Conversely, MDEO focuses on optimizing an identical task in multiple networks that may represent different complex systems. The goal of MDEO is to find the optimal solution for each domain.
Knowledge Transfer Across Networks: The solutions aiming at different tasks are aligned since MTEO is within the same domain. In the case of MDEO, the nodes or edges in one network are different from their counterparts in another network with the same index.
In this study, we take a further stride by delving into the case of more than three networks, i.e., many networks will be involved in the optimization. The proposed MDEO consists of four components 1) graph similarity 2) graph embedding 3) network alignment 4) many-network evolutionary optimization and its overview is as follows:
Multi-domain evolutionary optimization would present a significant challenge due to the inherent computational complexity when dealing with a large number of networks. Additionally, the efficacy of solutions on one network may not be necessarily translated to success on others. To address these issues, we propose a new measurement of graph similarity, quantifying the closeness between networks at the community level. The calculation of similarity enables us to focus on a subset of closely related networks during the knowledge transfer process, thereby effectively reducing the computational burden associated with multi-domain optimization. Moreover, we are able to prioritize the solutions from those networks with a higher degree of similarity based on the proposed graph similarity, thereby increasing the likelihood of successful adaptations.
In the multi-domain context, transferring solutions across different networks necessitates the establishment of node correspondences. We first employ a graph autoencoder to obtain graph embeddings aimed at capturing node similarity and higher-order interactions.
Building upon the derived graph embeddings, we then propose a novel network alignment method that combines supervised and unsupervised learning. In supervised learning, we propose a community-based anchor node selection method to build up the training set and improve the alignment accuracy. This approach empowers us to create mappings of nodes sharing analogous roles or structures across different networks.
In the process of many-network evolutionary optimization, we transfer solutions from highly similar networks to the target network, achieved by the utilization of the node mappings obtained through network alignment. By observing the contributions of these transferred solutions, we propose a self-adaptive model to determine the appropriate number of solutions to transfer between networks, with the consideration of the previously calculated graph similarity. Additionally, we develop a cross-domain mutation operator that makes better use of the knowledge from other networks to enhance the optimization.
In the validation with eight real-world networks, the experimental results reveal that the average fitness achieved within the MDEO framework outperforms that of classical evolutionary optimization, underscoring the effectiveness of our proposed framework. With the simulations of different attacks on the community, the results of metrics of normalized mutual information (NMI) [23] and adjusted rand index (ARI) [24] reveal that our proposed model can safeguard the community information from being exploited improperly.
In this work, we utilize the term ‘graph’ in the computational context and employ ‘network’ to denote real-world systems. Both terms refer to data composed of nodes and edges. The subsequent sections of the paper are organized as follows: Section II provides an overview of relevant literature, and Section III introduces the optimization problem–community deception. Section IV delves into the intricacies of our proposed MDEO framework, detailing its methodology and components. Section V entails an examination of the effectiveness of our method through a series of experiments conducted on various real-world networks. Lastly, Section VI concludes with a summary of our main results.
II Related works
The wide applicability of evolutionary algorithms in handling discrete problems with non-linear characteristics has led to their extensive utilization within complex networks. For example, Wang et al. [25] develop a multi-objective model by optimizing the computational complexity and nonuniform latency to enhance the tolerance of networks against attacks. Similarly, Wu et al. leverage the inherent community structure of networks to downsize the decision space and introduce a multi-objective framework for network reconstruction by considering the reconstruction error and sparsity in [26]. Based on the same optimization objectives, Ying et al. propose a logistic principal component analysis-based method to improve the representation of networks in [27].
Meanwhile, evolutionary algorithms have also been applied to address the problem of community deception. Chen et al. use the genetic algorithm to minimize the modularity after the modification on edges in [28]. Further, they categorize community deception based on the scales as community deception on node communities, target community and the entire community structure in [29]. A self-adaptive evolutionary framework is then proposed to facilitate the search of the optimal edge set and a permanence-based method is developed to reduce the search space in [30]. To alleviate the problem of the evaluation expense, Zhao and Cheong [31] propose a divide-and-conquer strategy to split the graph into subgraphs and optimize them with the genetic algorithm individually. Moreover, evolutionary algorithms have been applied to various other aspects of complex networks, including dynamic community detection [32], and module identification [33]. These applications demonstrate the versatility and effectiveness of evolutionary algorithms in tackling diverse challenges within the field of complex network analysis.
To date, considerable efforts have been directed towards tackling multi-task problems using evolutionary algorithms. In [5], Gupta et al. propose a paradigm, multi-factorial evolutionary algorithm, which can optimize different tasks simultaneously and has laid the foundation for the related research. For example, the optimization on single objective has been extended to the case of multiple objectives in [34], adopting the implicit transfer mode where the crossover is operated on the solutions aiming at different tasks [35, 36]. There is another route referred to explicit transfer that deliberately incorporates task-specific information to guide knowledge exchange between tasks by using statistic information [37] or constructing a mapping matrix [38, 39].
Beyond traditional fields, MTEO has been exploited within complex networks, such as in the optimization of network robustness. In [22], Wang et al. utilize the correlation of tasks to develop a new crossover operator, thereby obtaining knowledge from different individuals. As indicated in [40], structural destruction and cascading failure, two seemingly independent objectives, can promote each other in the optimization by developing a graph neural network-based method to exchange information between solutions aiming at different tasks. In one study of influence maximization [41], each task is treated as a transformation by formula, enabling the application of MTEO in this problem. In [21], Wu et al. promote the reconstruction of multiplex networks using MTEO with the inherent correlation of different layers. Similar work can be found in [42], in which the optimization of modularity in each layer is modeled as an individual task, thereby facilitating the community detection in multiplex networks.
Despite that networks often share the same characteristics, cross-domain evolutionary optimization, especially concerning network structures, has received limited attention. Motivated by this overlooked potential, we introduce a novel optimization framework—Multi-Domain Evolutionary Optimization.
III Optimization problem formulation
Community deception involves nodes intentionally concealing their true community affiliations by modifying (adding and removing) connections. In this study, we select this challenging combinatorial problem [43, 44] in complex networks as an illustration of our MDEO for the following concerns:
Exploiting Community Detection Algorithms: With the assistance of community detection algorithms, densely interconnected groups of nodes with high intra-group connectivity and low inter-group connectivity can be easily identified [42, 45]. The development of community detection, while useful for understanding network structures and interactions, can also have certain dangerous aspects caused by the potential misuse or exploitation of community information, as suggested in [46, 47].
Complexity of Community Deception: As an edge-level task, the search space of community deception is $|V|^{2}$, much higher than node-level tasks of $|V|$. Both existing and nonexistent edges will be considered in community deception, meaning the genes (edges) in the chromosomes (edge set) are heterogeneous. Therefore, community deception presents additional challenges than traditional tasks in complex networks such as the aforementioned critical node mining.
Community deception serves as an innovative advancement in network science, radically altering node affiliations through minimal topological adjustments. Specifically, the modification on the network involves the addition of nonexistent edges and the deletion of existing edges. Let $G=\{V,E\}$ be a network where $V$ and $E$ denote the nodes and edges, respectively. The updated edge set can be defined as follows:
$E^{\prime}=\left(E\cup E^{+}\right)\backslash E^{-},$ (1)
where $E^{+}$ and $E^{-}$ refers to the addition and deletion of edges, that is:
$\begin{array}[]{l}E^{+}\subseteq\{(u,v):u\in V\wedge v\in V,(u,v)\notin E\},\\ E^{-}\subseteq\{(u,v):u\in V\wedge v\in V,(u,v)\in E\},\end{array}$ (2)
with
$\left|E^{+}\right|+\left|E^{-}\right|\leq\beta,$ (3)
where $\beta$ refers to the perturbance budget, meaning the number of edges that can be modified.
Assume $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}^{\prime}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$ are the community partitions of original network $G$ and modified network $G^{\prime}$, then community deception can be represented as a combinatorial problem:
$\underset{\left\{E^{+},E^{-}\right\}}{\arg\max}\hskip 5.0pt\{\phi(\widetilde{\mathcal{C}},\widetilde{\mathcal{C}^{\prime})},E^{+},E^{-}\}.$ (4)
The function $\phi$ refers to the difference between the community structures, before and after modification.
IV Multi-Domain evolutionary optimization on community deception
In this section, the proposed MDEO will be introduced in detail, and its diagram can be found in Figure 1.
Figure 1: The diagram of multi-domain evolutionary optimization. Four networks representing different systems are optimized simultaneously. The transfer of knowledge across networks is achieved through the trained network alignment model. $e^{i}_{j}$ denote the $j$th edge in $i$th network and $M_{i,j}$ is the edge mapping from $i$th network to $j$th network. The solution with the colors of yellow and light red refers to the elite solution and transferred solution from other networks, respectively. The red edge is the mutated edge that will be replaced by the edge of elite solutions of other networks.
IV-A Graph similarity
The measurement of graph similarity involves the consideration of the interactions between nodes and higher-order structures, enabling us to perform selective knowledge transfer between networks. Instead of transferring all available information, it allows us to focus exclusively on pertinent domains, minimizing the risk of negative transfer and reducing the computational complexity.
In general, the graph similarity can be easily obtained byaveraging node similarity of networks [48] or computing the distance of graph embeddings [49, 50]. However, the former approach oversimplifies the complex interactions within the graph, ignoring the community structure and global patterns of networks. On the other hand, existing graph pooling techniques are typically trained on small graphs, potentially constraining their efficacy in accurately representing large-scale graphs.
In addition to overcoming the above problems, this similarity is also expected to facilitate the construction of training data for the network alignment model. To this end, we propose a community-based measurement of graph similarity. Given a network $G=(V,E)$ and its community distribution $\widetilde{\mathcal{C}}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\right\}$, for each community $\mathcal{C}\in\widetilde{\mathcal{C}}$, we can define $D_{\mathcal{C}}=\{d_{v}|v\in\mathcal{C}\}$ as the degree distribution of community $\mathcal{C}$ where $d_{v}$ is the node degree of $v$. With the consideration of the scales of different networks, we normalize this distribution by dividing the largest degree of the network, denoted as $D^{\prime}_{\mathcal{C}}=\{d_{v}/d^{max}|v\in\mathcal{C}\}$ where $d^{max}$ is the largest degree in the network $G$. In addition, another challenge emerges due to the varying lengths of these normalized distributions across networks, which complicates the direct comparison of community degree distributions. To address this issue, we introduce a new definition that unifies the degree distributions across communities of different scales into the same length.
Definition 1.
Degree interval-based distribution: Given a community $\mathcal{C}\in\widetilde{\mathcal{C}}$, then we can obtain the normalized degree distribution $D^{\prime}_{\mathcal{C}}=\left\{d^{\prime}_{1},d^{\prime}_{2},\ldots,d^{\prime}_{n}\right\}$ where $n$ is the number of nodes in community. Let $I_{k}=\left[a_{k},b_{k}\right)$ represent the $k$th interval, where $a_{k}$ is the lower bound, and $b_{k}$ is the upper bound of the interval, then we can have
$I_{j}=[(b_{j}-a_{j})*j,(b_{j}-a_{j})*(j+1)).$ (5)
For $x_{i}\in D^{\prime}_{\mathcal{C}}$, corresponding interval $I_{j}$ can be found such that: $x_{i}\in I_{j}$ where $j=\left\lfloor\frac{x_{i}}{b_{j}-a_{j}}\right\rfloor$. Then, we can calculate the number of data points $x_{i}\in D^{\prime}_{\mathcal{C}}$ that fall into each interval $I_{k}$: $h_{j}=\operatorname{Count}\left(x_{i}\right.$ such that $\left.x_{i}\in I_{j}\right)$ for each $j$. Here, we set $b_{j}-a_{j}=0.2$ for $j\in\{0,1,2,3,4\}$. The degree interval-based distribution can be obtained as follows:
$H_{\mathcal{C}}=\{\frac{h_{0}}{|\mathcal{C}|},\frac{h_{1}}{|\mathcal{C}|},\cdots,\frac{h_{l}}{|\mathcal{C}|}\},$ (6)
where $l=1/(b_{j}-a_{j})-1$.
With the above definition, we achieve a uniform representation of the degree distribution for each community, consistently comprising $l$ elements regardless of the number of nodes in the community.
Given two community structures $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2}\,\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2}\,\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, the similarity of communities $\mathcal{C}\in\widetilde{\mathcal{C}_{A}},\mathcal{C}^{\prime}\in\widetilde{\mathcal{C}_{B}}$ can be defined as follows:
$\operatorname{Diff}\left(\mathcal{C},\mathcal{C}^{\prime}\right)$ $=\frac{1}{2}\left(\mathbb{KL}\left(H_{\mathcal{C}}\|H_{\mathcal{C}^{\prime}}\right)+\mathbb{KL}\left(H_{\mathcal{C}^{\prime}}\|H_{\mathcal{C}}\right)\right)$
×$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|),$ (7)
where $\mathbb{KL}(.)$ refers to Kullback–Leibler divergence and the second item $\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|)$ is used to evaluate the difference of the scales of communities.
Traditional graph representations focus on pairwise relationships between nodes while high-order graphs extend this concept and it is able to capture interactions involving more than two nodes. Therefore, we also consider the high-order information of networks, allowing for a more refined comparison that captures structural patterns. Given a network $G$ and a motif $M$, the network $G$ can be converted into a weighted network $G_{M}$ as follows:
${G}^{{M}}=\left\{{V},{E}^{{M}}\right\},$ (8)
where ${E}^{M}\in E$ represents the set of edges within ${G}^{M}$. We can derive the adjacency matrix of hypergraph, denoted as $A_{M}$, in which each entry $(A_{M})_{ij}$ corresponds to the count of motifs that include the edge $e_{ij}$. We can measure the high-order difference between $\mathcal{C}$ and $\mathcal{C}^{\prime}$ as follows:
$\operatorname{Diff}^{M}\left(\mathcal{C},\mathcal{C}^{\prime}\right)$ $=\frac{1}{2}\left(\mathbb{KL}\left(H^{M}_{\mathcal{C}}\|H^{M}_{\mathcal{C}^{\prime}}\right)+\mathbb{KL}\left(H^{M}_{\mathcal{C}^{\prime}}\|H^{M}_{\mathcal{C}}\right)\right)$
×$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|).$ (9)
The similarity between communities is then defined as follows:
$\mathcal{S}(\mathcal{C},\mathcal{C}^{\prime})=\textbf{Exp}(-\operatorname{Diff}\left(\mathcal{C},\mathcal{C}^{\prime}\right))+\textbf{Exp}(-\operatorname{Diff}^{M}\left(\mathcal{C},\mathcal{C}^{\prime}\right))$ (10)
Definition 2.
Aligned community: Given two networks $G_{A}$, $G_{B}$ and their corresponding community structures $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, we can have the community-wise similarity matrix $\mathcal{S}\in\mathbb{R}^{k\times k^{\prime}}$ where $\mathcal{S}_{i,j}=\mathcal{S}(\mathcal{C}_{i},\mathcal{C}^{\prime}_{j}),i\in\{1,2,\cdots,k\},j\in\{1,2,\cdots,k^{\prime}\}$ Then, we define the set of aligned communities as follows:
$\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}=\{\{\mathcal{C}_{a_{1}},\mathcal{C}^{\prime}_{b_{1}}\},\{\mathcal{C}_{a_{2}},\mathcal{C}^{\prime}_{b_{2}}\},\cdots,\{\mathcal{C}_{a_{z}},\mathcal{C}^{\prime}_{b_{z}}\}\},$ (11)
where $z=\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)$. $\mathcal{C}_{a_{i}}$ and $\mathcal{C}^{\prime}_{b_{i}}$ refer to the $a_{i}$th and $b_{i}$th community in $\widetilde{\mathcal{C}_{A}}$ and $\widetilde{\mathcal{C}_{B}}$, respectively. Accordingly, we can have the set of their corresponding similarity values as follows:
$\mathcal{S}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}=\{\mathcal{S}_{a_{1},b_{1}},\mathcal{S}_{a_{2},b_{2}},\cdots,\mathcal{S}_{a_{z},b_{z}}\},$ (12)
which satisfies
$\mathcal{S}_{a_{i},b_{i}}=\mathbf{Max}(\mathcal{S}/\bigcup_{i>j}(\mathcal{S}_{a_{j},}\cup\mathcal{S}_{,b_{j}})).$ (13)
The similarity between $G_{A}$ and $G_{B}$ is defined as the average similarity of aligned community pairs.
$\mathcal{S}^{\mathcal{G}}_{A,B}=\frac{1}{\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)}\sum_{x\in\mathcal{S}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}}x.$ (14)
The process of calculating the graph similarity can be found in Algorithm 1.
Algorithm 1 Graph Similarity
Input: Networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$
Output: The quantitative similarity $S_{A,B}$ of $G_{A}$ and $G_{B}$.
1:Obtain high-order graphs ${G}^{m}_{A}=({V}_{A},{E}^{m}_{A})$ and ${G}^{m}_{B}=(V_{B},{E}^{{m}}_{B})$
2:Identify the community of $G_{A}$ and $G_{B}$ to obtain $\widetilde{\mathcal{C}}_{A}=\{\mathcal{C}_{1},\mathcal{C}_{2}\,\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}}_{B}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2}\,\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$
3:Compute the degree distribution of each community of $G$ and $G^{m}$
4:Normalize the degree distribution based on the scale of the network and transform it into the probability distribution
5:Construct the similarity matrix between the communities of the two networks based on K-L divergence
6:for $i=1$ to $\mathbf{Min}\{k,k^{\prime}\}$ do
7:     Find the largest value of $\mathcal{S}$ and its two corresponding communities as well as its index $i_{\text{min}},j_{\text{min}}=\text{argmin}S_{i,j}$ where $1\leq i\leq k,1\leq j\leq k^{\prime}$
8:     Update $\mathcal{S}^{\mathcal{G}}_{A,B}\leftarrow\mathcal{S}^{\mathcal{G}}_{A,B}+\mathbf{Max}(\mathcal{S})$
9:     Make the $i_{\text{min}}$ row and $j_{\text{min}}$ column of $S$ as 00
10:end for
11:Normalize the similarity $\mathcal{S}^{\mathcal{G}}_{A,B}$ based on the number of communities of $G_{A}$ and $G_{B}$
IV-B Graph embedding
To construct the node mappings across different networks, we employ graph autoencoders (GAE) to generate embeddings since it can capture meaningful and compact graph representations without the labeled data [51]. In GAE, the graph convolutional network (GCN) is taken as the encoder [52], i.e.,
$\mathbf{Z}=f_{encoder}(G)=\operatorname{GCN}(\mathbf{X},\mathbf{A}),$ (15)
where $\mathbf{A}$ denotes the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{N\times d}$ denotes nodes features. The information aggregation of GCN is defined as follows:
$\operatorname{GCN}(\mathbf{X},\mathbf{A})=\widetilde{\mathbf{A}}\operatorname{ReLU}\left(\widetilde{\mathbf{A}}\mathbf{X}\mathbf{W}_{\mathbf{0}}\right)\mathbf{W}_{\mathbf{1}},$ (16)
where $\mathbf{W}_{\mathbf{0}}$ and $\mathbf{W}_{\mathbf{1}}$ are learnable parameters and $\widetilde{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$.
The decoder takes the latent space representation $\mathbf{Z}$ as the input to reconstruct the adjacency matrix. The inner-product is adopted as the decoder function:
$\hat{\mathbf{A}}=f_{decoder}(\mathbf{Z})=\sigma\left(\mathbf{Z}\mathbf{Z}^{\mathrm{T}}\right),$ (17)
where $\sigma(\cdot)$ is the logistic sigmoid function.
GAE is trained by minimizing the reconstruction loss that measures the dissimilarity between the adjacency matrix $\mathbf{A}$ and the reconstructed matrix $\hat{\mathbf{A}}$. The loss function is:
$\mathcal{L}_{reconstruct}=-\frac{1}{N}\sum y\log\hat{y}+(1-y)\log(1-\hat{y}),$ (18)
where $y$ refers to the element in adjacency matrix $\mathbf{A}$ and $\hat{y}$ is the element that has the same position with $y$ in $\hat{\mathbf{A}}$.
IV-C Network alignment
To achieve knowledge transfer across different networks, we propose a new dual-learning network alignment model. It aims to find a correspondence (mapping) between nodes that have similar roles across different networks. The pseudocode of training network alignment is shown in Algorithm 2.
Definition 3.
Network alignment: Given two networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$, and their embeddings $\mathbb{E}_{A}$ and $\mathbb{E}_{B}$, network alignment model $\Phi$ is defined as a mapping from $G_{A}$ to $G_{B}$, such that for each $u\in V_{A}$ and its embedding $\mathbb{E}_{A}\left(u\right)$, we can have $\Phi^{a\rightarrow b}\left(\mathbb{E}_{A}\left(u\right)\right)=$ $\mathbb{E}_{B}\left(v\right),v\in V_{B}$ where $\{u,v\}$ is a pair of anchor nodes that play a similar role in their respective networks. We also denote the inverse mapping in a similar manner as $\Phi^{b\rightarrow a}\left(\mathbb{E}_{B}\left(v\right)\right)=\mathbb{E}_{A}\left(u\right)$.
In the supervised learning component, a set of anchor nodes is carefully selected in terms of the structural role in both networks, serving as labels to train the network alignment model. Given two networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$, we can then obtain their aligned community pairs $\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}$ as in Definition 2, which is the basis of the construction of anchor nodes. To make the network alignment model adapted to the cross-domain scenario, we consider both the important nodes and trivial nodes in each aligned community pair as anchor nodes.
Definition 4.
Anchor node (large degree) – $\mathcal{A}^{large}$: Let $v^{X,l}_{i}$ denote the $i$th largest-degree node in node set $X$. Then, given two networks $G_{A}$ and $G_{B}$, and an aligned community pair $\{\mathcal{C},\mathcal{C}^{\prime}\}\in\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}$, the set of anchor node of large degree is denoted as follows:
$\mathcal{A}^{large}=\bigcup\limits^{\{\mathcal{C},\mathcal{C}^{\prime}\}\in\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}\{v^{\mathcal{C},l}_{i},v^{\mathcal{C}^{\prime},l}_{i}\},i\in\{1,...,k_{l}\},$ (19)
where $k_{l}=\mathbf{Min}\{log_{2}|\mathcal{C}|,log_{2}|\mathcal{C}^{\prime}|\}$.
The loss for aligning the important nodes can be obtained as follows:
$\begin{split}\mathcal{L}_{large}&=\sum\limits_{\{u,v\}\in\mathcal{A}^{large}}dist(\Phi^{a\rightarrow b}(\mathbb{E}(u)),\mathbb{E}(v))\\ &\quad+dist(\Phi^{b\rightarrow a}(\mathbb{E}(v)),\mathbb{E}(u)).\end{split}$ (20)
where $dist$ refers to mean square error (MSE), measuring the difference between two embeddings.
To preserve the structural information, we will not consider the nodes with the least degree in the community. Instead, we consider the least-degree neighbors for the anchor nodes in $\mathcal{A}^{large}$. We thus have the following definition.
Definition 5.
Anchor node (small degree) – $\mathcal{A}^{small}$: Let $v^{X,s}_{i}$ denote the $i$th least-degree node in the node set $X$. Given an aligned pair of nodes $\{u,v\}\in\mathcal{A}^{large}$, we can define
$\mathcal{A}^{small}=\bigcup\limits^{\{u,v\}\in\mathcal{A}^{large}}\{v^{\mathcal{N}(u),s}_{i},v^{\mathcal{N}(v),s}_{j}\},i,j\in\{1,...,k_{s}\},$ (21)
where $k_{s}=\mathbf{Min}\{log_{2}|\mathcal{N}(u)|,log_{2}|\mathcal{N}(v)|\}$ and $\mathcal{N}(.)$ refers to the neighbor nodes.
These least important nodes are trivial in the network (and have little influence on the network structure), therefore, their roles are basically the same and will not require a one-to-one relationship. Thus, the loss function for small-degree nodes is based on the average of the similarity of their combination.
$\begin{split}\mathcal{L}_{small}&=\sum\limits_{\{u,v\}\in\mathcal{A}^{small}}\frac{1}{k_{s}}(dist(\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u)),\mathbb{E}_{B}(v))\\ &\quad+dist(\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v)),\mathbb{E}_{A}(u))).\end{split}$ (22)
Network alignment models $\Phi^{a\rightarrow b}$ and $\Phi^{b\rightarrow a}$ will also be trained in an unsupervised manner to enhance the mapping accuracy. Given a node $u\in V_{A}$, the automapping embedding $\Phi^{a\rightarrow b}(\Phi^{b\rightarrow a}(\mathbb{E}(u))$ should be as similar as the original embedding $\mathbb{E}(u)$, and the same for nodes $v\in V_{B}$. The automapping loss can be obtained as follows:
$\begin{split}\mathcal{L}_{us}&=dist(\Phi^{b\rightarrow a}(\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u)),\mathbb{E}_{A}(u))\\ &\quad+dist(\Phi^{a\rightarrow b}(\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v)),\mathbb{E}_{B}(v)).\end{split}$ (23)
Given two networks of $G_{A}$ and $G_{B}$, the loss function for training the mappings $\Phi^{a\rightarrow b}$ and $\Phi^{b\rightarrow a}$ is obtained as follows:
$\mathcal{L}_{alignment}\left(\mathbf{W}_{s,t},\mathbf{b}_{s,t},\mathbf{W}_{t,s},\mathbf{b}_{t,s}\right)=\mathcal{L}_{large}+\mathcal{L}_{small}+\mathcal{L}_{us}.$ (24)
Upon the optimization on $\mathcal{L}_{alignment}$, the trained mapping can align the remaining nodes between the two networks, enabling the solutions between networks to be transferred to each other.
Algorithm 2 Network Alignment
Input: Networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$
Output: Graph Mapping model $\mathbb{M}_{AB}$ and $\mathbb{M}_{BA}$
1:Obtain the embeddings of $V_{A}$ and $V_{B}$ as $\mathbb{E}_{A}=\mathbf{GAE}(G_{A})$ and $\mathbb{E}_{B}=\mathbf{GAE}(G_{B})$
2:for each epoch do
3:     #Unsupervised learning where $u\in V_{A}$ and $v\in V_{B}$
4:     Map the embedding $\mathbb{E}_{A}(u)$ to the space of $G_{B}$ via $\mathbb{M}_{AB}$, i.e., $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then map the embeddings of $G_{B}$ back to the space of $G_{A}$ via $\mathbb{M}_{BA}$
5:     Calculate the difference between $\Phi^{b\rightarrow a}(\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$ and $\mathbb{E}_{A}(u)$
6:     Map the embedding $\mathbb{E}_{B}(v)$ to the space of $G_{A}$ via $\mathbb{M}_{BA}$, i.e., $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$, and then map the embeddings of $G_{A}$ back to the space of $G_{B}$ via $\mathbb{M}_{AB}$
7:     Calculate the difference between $\Phi^{a\rightarrow b}(\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$ and $\mathbb{E}_{B}(v)$
8:     Calculate the unsupervised loss between the original embeddings and the automapped embeddings.
9:     #Supervised learning where $u\in V_{A}$ and $v\in V_{B}$, and (u,v) is an anchor pair
10:     Map $u$ to the space of $G_{B}$ via $\mathbb{M}_{AB}$ to obtain $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then calculating its difference with $\mathbb{E}_{A}(u)$
11:     Map $v$ to the space of $G_{A}$ via $\mathbb{M}_{BA}$ to obtain $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$ and then calculating its difference with $\mathbb{E}_{B}(v)$
12:     Use the anchor nodes to calculate the supervised loss
13:end for
14:Update the parameters of mappings $\mathbb{M}_{AB}$ and $\mathbb{M}_{BA}$
IV-D Many-network evolutionary optimization
In the evolutionary optimization of community deception, the chromosome is represented by the combination of genes (edges) with a parameter $\rho$ differentiating the edges to add or delete. The population of network $G_{i}\in\mathcal{G}$ can be defined as follows:
$P_{i}=\left[\begin{array}[]{cccccc}e_{1,1}^{i}&e_{1,2}^{i}&\cdots&\cdots&e_{1,\beta_{i}}^{i}&\rho_{1}^{i}\\ e_{2,1}^{i}&e_{2,2}^{i}&\cdots&\cdots&e_{2,\beta_{i}}^{i}&\rho_{2}^{i}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ e_{n,1}^{i}&e_{n,2}^{i}&\cdots&\cdots&e_{n,\beta_{i}}^{i}&\rho_{b}^{i}\end{array}\right],$ (25)
where the first $\rho$ edges are to add, and the rest $\beta-\rho$ are the edges to delete.
Let $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\ldots,\mathcal{C}^{\prime}_{k^{\prime}}\}$ represent two community partitions. The confusion matrix $m$ is employed to quantify the dissimilarity between these two community partitions. The entry $m_{ij}$ in the matrix denotes the number of common elements between $\mathcal{C}_{i}$ and $\mathcal{C}^{\prime}_{j}$. To facilitate the disengagement of nodes from their initial community affiliations, we can formulate
$F_{1}=-\sum_{i=1}^{|\widetilde{\mathcal{C}}|}\sum_{j=1}^{|\widetilde{\mathcal{C}}^{\prime}|}\frac{|\mathcal{C}_{i}|}{|V|}\left(\frac{m_{ij}}{|\mathcal{C}_{i}|}\log_{2}\frac{m_{ij}}{|\mathcal{C}_{i}|}\right).$ (26)
The objective of Eq. 26 is to foster decentralization within the community structure, albeit it may inadvertently result in numerous tiny communities. Ideally, the nodes from one community should integrate seamlessly into other communities without significantly altering the community sizes. To mitigate this extreme scenario, we introduce a robust alternative function that recalibrates the confusion matrix as follows:
$m^{\prime}_{ij}=\frac{m_{ij}}{|\mathcal{C}^{{}^{\prime}}_{j}|}*\log_{2}|\mathcal{C}^{{}^{\prime}}_{j}|.$ (27)
The first item, ${m_{ij}}/{|\mathcal{C}^{{}^{\prime}}_{j}|}$ refers to the proportion of the nodes of the original community $\mathcal{C}_{i}$ accounting for the new community $\mathcal{C}^{{}^{\prime}}_{j}$. The second term is utilized to prevent situations where a single node forms a community independently. Instead of resorting to computing a global indicator like entropy, we opt for an elegant yet effective approach by constraining its maximum value as follows:
$F_{2}=e^{-max(m^{\prime}_{ij})}.$ (28)
Then, the fitness function $F$ can be defined as follows:
$F=F_{1}*F_{2}.$ (29)
[…]
$\begin{cases}|\mathcal{G}^{\mathcal{S}}_{i}|=\sqrt{|\mathcal{G}|},\\ \forall G_{j}\in\mathcal{G}^{\mathcal{S}}_{i},\forall G_{k}\notin\mathcal{G}^{\mathcal{S}}_{i}:\mathcal{S}^{\mathcal{G}}_{i,j}>\mathcal{S}^{\mathcal{G}}_{i,k}.\end{cases}$ (30)
The transfer solution requires several rounds of reproduction to be integrated into the population of target domain. Combined with the interest of efficiency, we thus take $k=5$ empirically as the interval to determine whether to transfer or not. Let the best fitness values on $G_{i}$ in the current generation and $k$ generations ago be $F^{i}_{n\_0}$ and $F^{i}_{n\_k}$, respectively. Then, we can obtain the fitness difference between the current generation and $k$ generations ago and the fitness difference between $k$ and $2k$ generations ago as follows:
$d_{1}^{i}=\left|F_{n\_0}^{i}-F_{n\_k}^{i}\right|,\quad d_{2}^{i}=\left|F_{n\_k}^{i}-F_{n\_2k}^{i}\right|.$ (31)
Then, we define the transfer condition as follows:
$\begin{cases}d_{1,n}^{i}<d_{2,n}^{i},\\ n\geq 2k,\\ n\bmod k=0.\end{cases}$ (32)
The motivation of Eq.32 is that if the evolution speed is slower than the previous, then excellent solutions of other similar domains will be transferred in to facilitate the optimization.
To better use the transfer budget, we propose a self-adaptive mechanism to determine the number of solutions to transfer from $G\in\mathcal{G}^{\mathcal{S}}_{i}$ to $G_{i}$. Each transfer will be evaluated by its contribution to the new population. We will observe the difference between the current elite population $P^{Elite}_{i,n\_0}$ and the elite population before the last transfer $P^{Elite}_{i,n\_l}$, and its intersection with the last transfer ($P^{Transfer}_{i\leftarrow j,n\_l}$) from $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$. That is,
$I_{i\leftarrow j}^{\mathcal{G}}=\frac{|(P^{Elite}_{i,n\_0}\setminus P^{Elite}_{i,n\_l})\bigcap P^{Transfer}_{i\leftarrow j,n\_l}|}{P^{Elite}_{i,n\_0}}.$ (33)
The similarity matrix is then updated as follows:
$\mathcal{S}^{\mathcal{G}}_{i,j}\leftarrow I_{i\leftarrow j}^{\mathcal{G}}+\mathcal{S}^{\mathcal{G}}_{i,j},$ (34)
where $\mathcal{S}^{\mathcal{G}}_{i,j}$ is the similarity between $G_{i}$ and $G_{j}$.
In this manner, we are able to control the ratio of transferred solutions from different networks, given a fixed budget of transfer $\beta^{Transfer}$. The networks that are more similar to the target network should transfer more solutions, i.e.,
$|P^{Transfer}_{i\leftarrow j}|=\beta^{Transfer}*\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (35)
To better use the knowledge of other domains, we design a new mutation operator, which utilizes the elite edge of other networks to replace the edge in its own domain. Given a network $G_{i}\in\mathcal{G}$, the probability of the network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$ being selected to transfer is defined as follows:
$|P^{M}_{i\leftarrow j}|=\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (36)
Note that as different networks have different budgets $\beta$, the repairs to transferred solutions are necessary to ensure compatibility when the sizes of the transferred solution and the solution of the target network are not matched. Given an assisted network $G_{A}$ and a target network $G_{T}$, if $\beta_{A}>\beta_{T}$, then the edges in the transferred solution will be removed randomly until the length of the transferred solution meets the requirement. Otherwise, the edges will be sampled from existing edges as deletion and nonexistent edges as addition to make the transferred solution valid.
In the process of initialization, the chromosomes, denoting edge sets, undergo initialization through a sampling process, where both existing and nonexistent edges are included, with a random ratio. In selection, the roulette wheel strategy is utilized to select the chromosomes to participate in evolution. The chromosome with a larger fitness will be more likely to be selected. The elitism strategy is also adopted to safeguard against the interference of inferior genes, in which the top 10% of chromosomes with superior fitness are retained and directly propagated to the succeeding generation. Note that the operation of crossover and mutation may change the ratio of the added edge and deleted edge in the chromosome (edge set), thus the parameter $\rho$ will be updated during evolution.
Algorithm 3 Many-Network Evolutionary Optimization
Input: Networks $G_{i}=(V_{i},E_{i})\in\mathcal{G}$
Output: The optimized solution for the task on each network
1:Initialize population for each network
2:Initialize correlation coefficient matrix based on similarity matrix
3:Identify the top-k similar networks for each network
4:while the iteration does not reach the limit do
5:     for network $G_{i}\in\mathcal{G}$ do
6:         Perform selection
7:         if condition meets then
8:              for network $G_{j}$ in $\mathcal{G}^{\mathcal{S}}_{i}$ do
9:                  Compute the number of solutions transferred from $G_{j}$
10:                  Transfer elitism solutions from $G_{j}$ and add them to the population $P_{i}$
11:              end for
12:              Update and normalize correlation matrix
13:         end if
14:         Perform crossover operation
15:         Perform mutation operation
16:     end for
17:end while
V Experimental studies
In this section, eight real-world networks are utilized to examine our proposed MDEO, and the parameter sensitivity and structural change are also investigated.
V-A Experimental setting
We use eight different real-world networks [53] of various sizes from different domains to validate our proposed MDEO. In the process of evolution, the population size of each network is set to 100. In the reproduction, the probabilities of crossover and mutation are set to 0.5 and 0.1, respectively. In the comparison, the effectiveness of random modification will be investigated and the pioneering work in community deception, DICE [43] will also used as one of baselines. We also optimize the modularity that is the most recognized metric in the studies of community, using the genetic algorithm as in [28] to conceal community information. To demonstrate the effectiveness, the classical evolutionary optimization, referred to single-domain evolutionary optimization (SDEO), will also be included in the comparison.
V-B Experiment on real-world networks
In Table II(b), we compare the efficacy of different community deception algorithms by examining two metrics of NMI and ARI. The lower NMI and ARI are, the better protection the community deception algorithm achieves. We also use Wilcoxon rank sum tests to observe the statistical difference between the compared methods and MDEO, with the significance level $p=0.05$. The symbols ‘+’ and ‘-’ means the compared method is better/worse than MDEO and $\approx$’ means that the compared method has the similar performance with MDEO. The community deception algorithms are tested with three attackers (community detection algorithms) on eight networks of different domains. NMI and ARI basically exhibit the same trend but there are still some exceptions. For example, the NMI value of SDEO and MDEO is higher than DICE while their ARI values are much lower than that of DICE. As observed, our proposed MDEO clearly outperforms other methods in most cases and is better than SDEO, suggested by its lower average NMI and ARI than other compared methods. This result indicates the effectiveness of our method in securing community information. On the other hand, the result is subject to the attacker. For example, the NMI and ARI in Adjnoun when dealing with the attack from FastGreedy, are lower than other networks. Network is also an important variable, as suggested by that NMI and ARI in Adjnoun are higher when dealing with WalkTrap than other networks.
TABLE I: The comparison of the mean and standard deviation of NMI and ARI obtained from different community deception algorithms. The results are averaged from 20 independent runs. Three community detection algorithms are tested, i.e., FastGreedy and WalkTrap.
Network NMI ARI
RAM DICE Mod SDEO MDEO RAM DICE Mod SDEO MDEO
Dolphins 0.75$\pm$0.060(-) 0.74$\pm$0.082(-) 0.76$\pm$0.094(-) 0.51$\pm$0.056(-) 0.43$\pm$0.041 0.70$\pm$0.072(-) 0.69$\pm$0.091(-) 0.71$\pm$0.115(-) 0.31$\pm$0.053(-) 0.24$\pm$0.032
Lesmis 0.79$\pm$0.113(-) 0.75$\pm$0.118(-) 0.80$\pm$0.093(-) 0.55$\pm$0.024(-) 0.47$\pm$0.035 0.72$\pm$0.178(-) 0.65$\pm$0.187(-) 0.76$\pm$0.113(-) 0.31$\pm$0.026(-) 0.25$\pm$0.025
Polbooks 0.85$\pm$0.100(-) 0.89$\pm$0.079(-) 0.92$\pm$0.060(-) 0.54$\pm$0.047($\approx$) 0.52$\pm$0.035 0.84$\pm$0.117(-) 0.89$\pm$0.093(-) 0.92$\pm$0.079(-) 0.45$\pm$0.072(-) 0.41$\pm$0.054
Adjnoun 0.54$\pm$0.118(-) 0.42$\pm$0.094(-) 0.48$\pm$0.118(-) 0.24$\pm$0.036(-) 0.24$\pm$0.026 0.43$\pm$0.159(-) 0.29$\pm$0.119(-) 0.37$\pm$0.154(-) 0.09$\pm$0.018(-) 0.07$\pm$0.022
Erdos 0.68$\pm$0.090(-) 0.67$\pm$0.067(-) 0.70$\pm$0.080(-) 0.50$\pm$0.028(+) 0.49$\pm$0.017 0.58$\pm$0.134(-) 0.58$\pm$0.103(-) 0.61$\pm$0.125(-) 0.28$\pm$0.026(+) 0.27$\pm$0.016
USAir 0.73$\pm$0.114(-) 0.65$\pm$0.126(-) 0.73$\pm$0.111(-) 0.48$\pm$0.057(-) 0.41$\pm$0.045 0.70$\pm$0.168(-) 0.64$\pm$0.163(-) 0.74$\pm$0.139(-) 0.34$\pm$0.055(-) 0.27$\pm$0.050
Netscience 0.95$\pm$0.029(-) 0.93$\pm$0.032(-) 0.96$\pm$0.026(-) 0.83$\pm$0.015(-) 0.80$\pm$0.020 0.90$\pm$0.071(-) 0.85$\pm$0.085(-) 0.91$\pm$0.068(-) 0.57$\pm$0.029(-) 0.53$\pm$0.036
BioCelegans 0.52$\pm$0.058(-) 0.49$\pm$0.042(-) 0.49$\pm$0.067(-) 0.40$\pm$0.036($\approx$) 0.39$\pm$0.029 0.42$\pm$0.084(-) 0.37$\pm$0.073(-) 0.39$\pm$0.010(-) 0.22$\pm$0.033($\approx$) 0.21$\pm$0.024
(a) FastGreedy
Network NMI ARI
RAM DICE Mod SDEO MDEO RAM DICE Mod SDEO MDEO
Dolphins 0.67$\pm$0.063(-) 0.64$\pm$0.077(-) 0.67$\pm$0.055(-) 0.57$\pm$0.031(-) 0.54$\pm$0.037 0.51$\pm$0.092(-) 0.48$\pm$0.088(-) 0.51$\pm$0.077(-) 0.30$\pm$0.029(-) 0.27$\pm$0.034
Lesmis 0.90$\pm$0.039(-) 0.86$\pm$0.047(-) 0.90$\pm$0.044(-) 0.81$\pm$0.022(-) 0.74$\pm$0.024 0.80$\pm$0.101(-) 0.72$\pm$0.109(-) 0.77$\pm$0.120(-) 0.55$\pm$0.027(-) 0.48$\pm$0.023
Polbooks 0.92$\pm$0.060(-) 0.88$\pm$0.044(-) 0.90$\pm$0.072(-) 0.64$\pm$0.017($\approx$) 0.63$\pm$0.014 0.93$\pm$0.064(-) 0.88$\pm$0.056(-) 0.89$\pm$0.122(-) 0.49$\pm$0.021($\approx$) 0.48$\pm$0.017
Adjnoun 0.79$\pm$0.035(-) 0.76$\pm$0.051(-) 0.80$\pm$0.042(-) 0.72$\pm$0.022(+) 0.72$\pm$0.018 0.52$\pm$0.111(-) 0.48$\pm$0.150(-) 0.52$\pm$0.126(-) 0.20$\pm$0.035(+) 0.20$\pm$0.022
Erdos 0.90$\pm$0.014(-) 0.87$\pm$0.022(-) 0.88$\pm$0.021(-) 0.85$\pm$0.012(+) 0.85$\pm$0.010 0.70$\pm$0.055(-) 0.62$\pm$0.071(-) 0.63$\pm$0.074(-) 0.47$\pm$0.036($\approx$) 0.45$\pm$0.031
USAir 0.83$\pm$0.045(-) 0.82$\pm$0.066(-) 0.83$\pm$0.041(-) 0.77$\pm$0.014(-) 0.75$\pm$0.009 0.77$\pm$0.070(-) 0.74$\pm$0.139(-) 0.76$\pm$0.103(-) 0.48$\pm$0.044(-) 0.44$\pm$0.036
Netscience 0.94$\pm$0.027(-) 0.92$\pm$0.023(-) 0.93$\pm$0.024(-) 0.86$\pm$0.010(-) 0.85$\pm$0.006 0.82$\pm$0.108(-) 0.75$\pm$0.102(-) 0.80$\pm$0.095(-) 0.53$\pm$0.026(-) 0.51$\pm$0.017
BioCelegans 0.78$\pm$0.029(-) 0.72$\pm$0.028(-) 0.79$\pm$0.030(-) 0.74$\pm$0.011(-) 0.73$\pm$0.009 0.45$\pm$0.060(-) 0.50$\pm$0.063(-) 0.47$\pm$0.081(-) 0.31$\pm$0.017(-) 0.30$\pm$0.017
(b) WalkTrap
To examine the effectiveness of our proposed MDEO, we compare it with the traditional optimization SDEO by observing their evolution process. As observed in Figures 3a and 3b, SDEO converges around or before $50th$ generation while MDEO still exhibits an upward trend even in $200th$ generation on all networks, indicating that MDEO can effectively enhance the optimization and is not easily trapped into local optimality. It is worth noting that the optimization curves are ladder-shaped, meaning the solution transferred from other similar networks can improve the optimization on the target network. All curves depict an upward trend as shown in Figures 3a and 3b, depicting the high robustness of MDEO.
* Refer to caption
Figure 2: The evolution process of MDEO and SDEO. The attacker is FastGreedy.
Figure 3: The evolution process of MDEO and SDEO. The attacker is WalkTrap.
V-C Parameter sensitivity analysis
To bolster the completeness of our study, we conduct a targeted sensitivity analysis. In order to validate whether the transfer contributes to the optimization, we investigate the influence of the number of networks assisting in the target network on the results, as shown in Table II. For fairness, the number of solutions to transfer is the same for each setting. As seen, the results are related to the number of assisted networks, as reflected by that the fitness values of $|\mathcal{G}^{\mathcal{S}}|=2$ and $|\mathcal{G}^{\mathcal{S}}|=3$ are larger than that of $|\mathcal{G}^{\mathcal{S}}|=1$. Furthermore, in the test of a total of eight networks, $|\mathcal{G}^{\mathcal{S}}|=3$ achieves the optimum on five networks, five networks, and seven networks respectively when defending against FastGreedy and WalkTrap, meaning our transfer strategy is effective in the optimization.
We also examine the sensitivity of the ratio of elite solutions that will be directly transferred into the next generation in evolution. We refer this ratio as $\mu$ and set it to $\mu_{1}=0.05$, $\mu_{2}=0.10$, $\mu_{3}=0.15$ and $\mu_{4}=0.20$, respectively. We can observe that no ratio exhibits obvious advantages over other settings. For example, the large ratio of elite $\mu_{4}$ (purple line) achieves the best performance on Adjnoun, Dolphins, Lesmis, Erods and Polbooks while achieving the worst performance on USAir and BioCelegans. On the contrary, the setting of low ratio $\mu_{1}$, which has the best performance on USAir and BioCelegans, is inferior to other settings on the rest of networks. These two examples demonstrate that MDEO is not sensitive to the ratio of elite solutions.
TABLE II: The fitness comparison of the number of networks participating in transfer in the process of optimization.
Network FastGreedy WalkTrap
$|\mathcal{G}^{\mathcal{S}}|=1$ $|\mathcal{G}^{\mathcal{S}}|=2$ $|\mathcal{G}^{\mathcal{S}}|=3$ $|\mathcal{G}^{\mathcal{S}}|=1$ $|\mathcal{G}^{\mathcal{S}}|=2$ $|\mathcal{G}^{\mathcal{S}}|=3$
Dolphins 0.177 0.198 0.199 0.213 0.231 0.237
Lesmis 0.190 0.193 0.195 0.110 0.105 0.113
Polbooks 0.044 0.051 0.049 0.050 0.051 0.047
Adjnoun 0.684 0.653 0.649 0.302 0.297 0.312
Erdos 0.141 0.156 0.159 0.046 0.060 0.046
USAir 0.031 0.038 0.040 0.023 0.022 0.024
Netscience 0.031 0.038 0.039 0.041 0.040 0.052
BioCelegans 0.101 0.110 0.106 0.086 0.094 0.100
Figure 4: The fitness profile comparison of different elite transfer ratios.
V-D Structural change
The structural change in some indices after perturbation is shown in Table III, through which we can observe which kind of edges that the community deception algorithm prefers to modify. CC refers to the clustering coefficient, measuring the extent to which nodes tend to cluster together. The clustering coefficients of all networks have decreased, meaning the local connectivity within the network is becoming lower. On the other hand, the average shortest distance (ASD) gets improved after rewiring edges, which suggests the distance between communities is shortened. As for the centralities, the change in Betweenness is universally larger than the change in PageRank, indicating the proposed community deception algorithm tends to change the bridge edges. The observed reduction in modularity indicates that our algorithm effectively weakens the internal connections of individual communities while simultaneously enhancing their external connectivity.
TABLE III: The structural change after modifying edges obtained by MDEO-II. The tested attacker is FastGreedy.
Data $|E^{+}\cup E^{-}|$ CC ASD $20\%-$ Betweenness $20\%-$ PageRank Modularity
Dolphins 10 $0.309\rightarrow 0.291$ $3.357\rightarrow 3.051$ $1\rightarrow 0.667$ $1\rightarrow 0.933$ $0.495\rightarrow 0.442$
Lesmis 10 $0.498\rightarrow 0.490$ $2.641\rightarrow 2.582$ $1\rightarrow 0.913$ $1\rightarrow 0.916$ $0.500\rightarrow 0.489$
Polbooks 20 $0.348\rightarrow 0.334$ $3.079\rightarrow 2.830$ $1\rightarrow 0.702$ $1\rightarrow 0.974$ $0.502\rightarrow 0.472$
Adjnoun 20 $0.157\rightarrow 0.156$ $2.536\rightarrow 2.523$ $1\rightarrow 0.923$ $1\rightarrow 0.975$ $0.294\rightarrow 0.272$
Netscience 30 $0.431\rightarrow 0.413$ $6.042\rightarrow 5.322$ $1\rightarrow 0.844$ $1\rightarrow 0.946$ $0.839\rightarrow 0.798$
Erdos 50 $0.214\rightarrow 0.206$ $4.021\rightarrow 3.958$ $1\rightarrow 0.932$ $1\rightarrow 0.958$ $0.513\rightarrow 0.484$
BioCelegans 50 $0.124\rightarrow 0.123$ $2.664\rightarrow 2.664$ $1\rightarrow 0.948$ $1\rightarrow 0.959$ $0.395\rightarrow 0.390$
USAir 100 $0.396\rightarrow 0.386$ $2.738\rightarrow 2.715$ $1\rightarrow 0.902$ $1\rightarrow 0.973$ $0.319\rightarrow 0.303$
VI Conclusion
In this work, we have explored a framework–multi-domain evolutionary optimization. We actualized this concept within the realm of network structures, predicated on the inherent shared properties of real-world networks. The proposed MDEO consists of three components: 1) measuring the similarity of domains at the community level; 2) selecting the anchor pair nodes and training the mapping of solutions of different domains 3) optimizing different domains simultaneously and making them promote each other. The proposed method was validated on the task of community deception with eight different real-world networks of varying sizes from various domains, and the experimental results (with statistical analysis and sensitivity analysis) show the fitness values of evolutionary optimization on different networks have been improved, suggesting the effectiveness of our proposed multi-domain evolutionary optimization. As an exploratory work, MDEO has been successfully applied to networks of similar sizes. The knowledge transfer from a large network to a small network or vice versa, has not been studied yet. In future work, we will develop a more robust and adaptive framework that allows networks of various scales to be optimized simultaneously with improved effectiveness.

### v2

@@
-Multi-Domain Evolutionary Optimization of Network Structures
+Multi-Domain Evolutionary Optimization on Combinatorial Problems in Complex Networks
 Abstract
-Multi-Task Evolutionary Optimization (MTEO), an important field focusing on addressing complex problems through optimizing multiple tasks simultaneously, has attracted much attention.
-While MTEO has been primarily focusing on task similarity, there remains a hugely untapped potential in harnessing the shared characteristics between different domains to enhance evolutionary optimization.
+Knowledge transfer-based evolutionary optimization has garnered significant attention, such as in multi-task evolutionary optimization (MTEO), which aims to solve complex problems by simultaneously optimizing multiple tasks.
+While this emerging paradigm has been primarily focusing on task similarity, there remains a hugely untapped potential in harnessing the shared characteristics between different domains.
 For example, real-world complex systems usually share the same characteristics, such as the power-law rule, small-world property and community structure, thus making it possible to transfer solutions optimized in one system to another to facilitate the optimization.
-Drawing inspiration from this observation of shared characteristics within complex systems, we set out to extend MTEO to a novel framework—multi-domain evolutionary optimization (MDEO).
-To examine the performance of the proposed MDEO, we utilize a challenging combinatorial problem of great security concern—community deception in complex networks as the optimization task.
-To achieve MDEO, we propose a community-based measurement of graph similarity to manage the knowledge transfer among domains.
-Furthermore, we develop a graph representation-based network alignment model that serves as the conduit for effectively transferring solutions between different domains.
-Moreover, we devise a self-adaptive mechanism to determine the number of transferred solutions from different domains and introduce a novel mutation operator based on the learned mapping to facilitate the utilization of knowledge from other domains.
-Experiments on eight real-world networks of different domains demonstrate MDEO’s superiority in efficacy compared to classical evolutionary optimization.
-Simulations of attacks on the community validate the effectiveness of the proposed MDEO in safeguarding community security.
+Drawing inspiration from this observation of shared characteristics within complex systems, we present a novel framework, multi-domain evolutionary optimization (MDEO).
+First, we propose a community-level measurement of graph similarity to manage the knowledge transfer among domains.
+Furthermore, we develop a graph learning-based network alignment model that serves as the conduit for effectively transferring solutions between different domains.
+Moreover, we devise a self-adaptive mechanism to determine the number of transferred solutions from different domains, and introduce a knowledge-guided mutation mechanism that adaptively redefines mutation candidates to facilitate the utilization of knowledge from other domains.
+To evaluate its performance, we use a challenging combinatorial problem known as adversarial link perturbation as the primary illustrative optimization task.
+Experiments on multiple real-world networks of different domains demonstrate the superiority of the proposed framework in efficacy compared to classical evolutionary optimization.
 Index Terms:
-Complex network, multi-domain evolutionary optimization, multi-task evolutionary optimization, transfer learning.
+Complex network, multi-domain evolutionary optimization, knowledge transfer, combinatorial problem.
 I Introduction
-Multi-task evolutionary optimization (MTEO) has stood as an emerging framework to solve complex optimization problems, undergoing remarkable advancements [1, 2, 3, 4] since the multi-factorial evolutionary algorithm (MFEA) was proposed [5].
-MTEO aims to address situations where there are multiple related tasks or objectives being optimized simultaneously, with the goal of achieving better performance on each task.
-While most existing endeavors primarily concentrate on exploring relationships between tasks, a research gap persists regarding correlations between different domains in evolutionary optimization.
-In real-world applications, we encounter a multitude of networks that capture different aspects of complex systems [6].
-These networks may represent social networks [7] or transportation networks [8].
-Many of these networks share common structural properties, such as power-law distribution [9, 10], community structure [11, 12] and small-world characteristic [13, 14, 15].
-In real-world scenarios, different domains often have common objective, such as the task of critical node mining in complex networks [16, 17], which can contribute to the prevention of catastrophic outages in the power networks [18], the identification of drug target candidates in protein networks [19], or the improvement of the robustness in communication networks [20].
-The existence of shared attributes and structural properties across real-world networks presents an avenue for knowledge transfer and collaborative learning when optimizing an identical task.
-To date, multi-task evolutionary optimization of complex networks has already been extensively studied [21, 22], however, leveraging the correlation of domains to facilitate the optimization of the network structure remains underexplored.
-Inspired by MTEO, we propose a new framework called multi-domain evolutionary optimization (MDEO) for network structures, in which solutions obtained from optimizing one network can be effectively transferred and adapted to optimize another network with improved efficacy.
+Knowledge transfer in evolutionary optimization has emerged as a promising paradigm for tackling complex optimization problems by enabling the sharing of useful information across related problems.
+One representative framework is multi-task evolutionary optimization (MTEO), which has seen notable progress in recent years [1, 2], following the introduction of the multi-factorial evolutionary algorithm (MFEA) [3].
+MTEO is designed to handle scenarios involving multiple related tasks being optimized concurrently, with the aim of leveraging inter-task similarity to improve performance on each task [4, 5].
+I-A Motivation
+In this work, we shift the focus from task-based settings to the structural properties of the network: While most existing endeavors primarily concentrate on exploring relationships between tasks, a research gap persists regarding correlations between different domains in evolutionary optimization.
+In real-world applications, we encounter a multitude of networks representing different complex systems [6].
+These networks may represent social networks [7] or transportation networks [8], many of which share common structural properties, such as power-law distribution [9, 10], community structure [11, 12] and small-world characteristic [13, 14, 15].
+In real-world scenarios, the same objective often exists in different domains, such as the task of critical node mining in complex networks [7, 16, 17], which can contribute to the prevention of catastrophic outages in power networks [18], the identification of drug target candidates in protein networks [19], or the improvement of the robustness in communication networks [20].
+The existence of shared attributes and structural properties across real-world networks presents an avenue for knowledge transfer and collaborative optimization when dealing with an identical task across different systems.
+To date, MTEO in complex networks has already been extensively studied [21, 22], however, leveraging the correlation of domains to facilitate the optimization of network structure remains underexplored.
+Therefore, we propose a new framework called multi-domain evolutionary optimization (MDEO) for network-structured combinatorial problems, in which solutions obtained from optimizing one network can be effectively transferred and adapted to optimize another network with improved efficacy.
 The following are the main differences between MTEO and MDEO, justifying the novelty of our work:
-Granularity of Tasks: In MTEO, the emphasis is on optimizing different tasks, which may be distinct and well-defined within a single network.
-Conversely, MDEO focuses on optimizing an identical task in multiple networks that may represent different complex systems.
+Granularity of Tasks: In MTEO, the emphasis is on optimizing different tasks within a single domain.
+Conversely, MDEO focuses on optimizing an identical task in multiple networks representing different complex systems.
 The goal of MDEO is to find the optimal solution for each domain.
-Knowledge Transfer Across Networks: The solutions aiming at different tasks are aligned since MTEO is within the same domain.
-In the case of MDEO, the nodes or edges in one network are different from their counterparts in another network with the same index.
-In this study, we take a further stride by delving into the case of more than three networks, i.e., many networks will be involved in the optimization.
-The proposed MDEO consists of four components 1) graph similarity 2) graph embedding 3) network alignment 4) many-network evolutionary optimization and its overview is as follows:
-Multi-domain evolutionary optimization would present a significant challenge due to the inherent computational complexity when dealing with a large number of networks.
-Additionally, the efficacy of solutions on one network may not be necessarily translated to success on others.
+Note that if multiple tasks concurrently exist across different domains, MDEO could be extended to multi-domain multi-task evolutionary optimization.
+Knowledge Transfer Across different domains: In MDEO, knowledge is transferred across distinct domains, such as the social network, power network and biology network, enhancing solutions by leveraging insights from one domain to another.
+This contrasts with MTEO, where knowledge is transferred within the same domain across tasks with similar characteristics, whether continuous or combinatorial.
+I-B Contributions
+In this study, we take a further stride and delve into the case of more than three networks, i.e., many networks will be involved in the optimization.
+The proposed MDEO consists of four components 1) graph similarity; 2) graph embedding; 3) network alignment; 4) many-network evolutionary optimization and its overview is as follows:
+Multi-domain evolutionary optimization may present a significant challenge due to the inherent computational complexity when dealing with a large number of networks.
+Additionally, the efficacy of solutions on one network may not necessarily translate to success on others.
 To address these issues, we propose a new measurement of graph similarity, quantifying the closeness between networks at the community level.
-The calculation of similarity enables us to focus on a subset of closely related networks during the knowledge transfer process, thereby effectively reducing the computational burden associated with multi-domain optimization.
-Moreover, we are able to prioritize the solutions from those networks with a higher degree of similarity based on the proposed graph similarity, thereby increasing the likelihood of successful adaptations.
+The similarity measurement enables us to focus on a subset of closely related networks during the knowledge transfer process, thereby effectively reducing the computational burden and increasing the likelihood of successful adaptations.
 In the multi-domain context, transferring solutions across different networks necessitates the establishment of node correspondences.
-We first employ a graph autoencoder to obtain graph embeddings aimed at capturing node similarity and higher-order interactions.
+We first employ the graph autoencoder to obtain graph embeddings aimed at capturing node similarity and higher-order interactions.
 Building upon the derived graph embeddings, we then propose a novel network alignment method that combines supervised and unsupervised learning.
-In supervised learning, we propose a community-based anchor node selection method to build up the training set and improve the alignment accuracy.
+In supervised learning, we propose a community-level anchor node selection method to build up the training set and improve the alignment accuracy.
 This approach empowers us to create mappings of nodes sharing analogous roles or structures across different networks.
-In the process of many-network evolutionary optimization, we transfer solutions from highly similar networks to the target network, achieved by the utilization of the node mappings obtained through network alignment.
-By observing the contributions of these transferred solutions, we propose a self-adaptive model to determine the appropriate number of solutions to transfer between networks, with the consideration of the previously calculated graph similarity.
-Additionally, we develop a cross-domain mutation operator that makes better use of the knowledge from other networks to enhance the optimization.
-In the validation with eight real-world networks, the experimental results reveal that the average fitness achieved within the MDEO framework outperforms that of classical evolutionary optimization, underscoring the effectiveness of our proposed framework.
-With the simulations of different attacks on the community, the results of metrics of normalized mutual information (NMI) [23] and adjusted rand index (ARI) [24] reveal that our proposed model can safeguard the community information from being exploited improperly.
-In this work, we utilize the term ‘graph’ in the computational context and employ ‘network’ to denote real-world systems.
-Both terms refer to data composed of nodes and edges.
-The subsequent sections of the paper are organized as follows: Section II provides an overview of relevant literature, and Section III introduces the optimization problem–community deception.
+In the process of many-network evolutionary optimization, we transfer solutions from similar networks to the target network, achieved by the node mappings obtained through network alignment.
+By observing the contributions of these transferred solutions, we leverage a self-adaptive model to determine the appropriate number of solutions to transfer between networks, with the consideration of the previously calculated graph similarity.
+Validation on eight real-world networks for edge-level tasks reveals that the MDEO framework achieves higher average fitness than classical evolutionary optimization.
+This trend is also observed in the node-level tasks, indicating the generalizability and adaptability of MDEO.
+I-C Organization
+The subsequent sections of the paper are organized as follows: Section II provides an overview of related work, and Section III introduces the optimization problem–community deception.
 Section IV delves into the intricacies of our proposed MDEO framework, detailing its methodology and components.
 Section V entails an examination of the effectiveness of our method through a series of experiments conducted on various real-world networks.
-Lastly, Section VI concludes with a summary of our main results.
-II Related works
-The wide applicability of evolutionary algorithms in handling discrete problems with non-linear characteristics has led to their extensive utilization within complex networks.
+Section VI discusses the possible extension of MDEO.
+Lastly, Section VII concludes with a summary of our main results.
+II Related work
+The capability of evolutionary optimization in handling discrete problems with non-linear characteristics has led to their extensive utilization within complex systems.
+For instance, Wang et al.
+[23] developed a multi-objective model by considering the nonuniform latency and computational complexity to enhance the tolerance of networks against attacks.
+Similarly, Wu et al.
+[24] leveraged the community distributions of networks to downsize the search space and introduce a multi-objective framework for network reconstruction by optimizing the reconstruction error and sparsity.
+Based on the same optimization objectives, Ying et al.
+[25] proposed a logistic principal component analysis-based method to improve the representation of networks.
+Moreover, evolutionary optimization has also been used in the diffusion source localization for sensor deployment to enhance the identification accuracy [26].
+Meanwhile, evolutionary algorithms were utilized to address the problem of community deception as well.
+Chen et al.
+[27] used the genetic algorithm to minimize the modularity and suggest the modifications to edges.
+Further, they categorized this problem into node communities, target communities, and all communities according to the scales [28].
+Subsequently, a self-adaptive evolutionary framework is designed to streamline the search for the optimal edge set, complemented by the development of a permanence-based method to minimize the search space [29].
+To mitigate the costs associated with evaluation, Zhao and Cheong [30] developed a divide-and-conquer strategy that partitions the network into sub-components, each optimized individually and cooperatively using the evolutionary algorithm.
+More applications of evolutionary algorithms on complex systems include dynamic community detection [31], and module identification [32].
+These applications demonstrate the versatility and effectiveness of evolutionary algorithms in tackling diverse challenges within the field of complex network analysis.
+Recently, considerable efforts have been directed towards tackling evolutionary multi-task cases where the development of multi-factorial evolutionary algorithm [3] has laid the foundation for related studies.
+In addition, the optimization on single objective has been extended to the case of multiple objectives [33], adopting the implicit transfer mode where the crossover is operated on the solutions aiming at different tasks [34, 35].
+In another extension of MTEO to the multi-objective setting [36], each task is decomposed into multiple subproblems, and the transfer probability of a solution is assessed based on how much it improves the performance of its corresponding subproblem.
+There is another route referred to explicit transfer that uses task-specific information to guide knowledge exchange between tasks by using statistical information [37] or constructing a mapping matrix [38, 39].
+The primary focus of MTEO research lies in enhancing transfer strategies.
+For example, Wu et al.
+[40] proposed a diversified reasoning approach that expands solution space via varied transfer patterns.
+Xue et al.
+[41] used neural networks to evaluate task similarity and guide information flow, thereby enhancing transfer quality.
+In [42], reinforcement learning is used to learn effective policies where the knowledge transfer is modeled as sequential decisions.
+TABLE I: Comparison of evolutionary optimization frameworks based on knowledge transfer.
+Multi-task evolutionary optimization is divided into two categories: MTEO-ConO, which addresses continuous problems, and MTEO-ComO, which focuses on combinatorial problems.
+  Framework Domain Scope Problem Type Space Type Task Scope
+MTEO-ConO [3, 33, 34, 35, 43, 44] Single-domain Continuous optimization Continuous Multiple
+MTEO-ComO [21, 38, 45, 46, 47, 48, 49] Single-domain Combinatorial optimization Discrete Multiple
+MDEO Multi-domain Combinatorial optimization Discrete Single/Multiple
+Beyond conventional areas, MTEO has been leveraged to strengthen the robustness of complex networks.
 For example, Wang et al.
-[25] develop a multi-objective model by optimizing the computational complexity and nonuniform latency to enhance the tolerance of networks against attacks.
-Similarly, Wu et al.
-leverage the inherent community structure of networks to downsize the decision space and introduce a multi-objective framework for network reconstruction by considering the reconstruction error and sparsity in [26].
-Based on the same optimization objectives, Ying et al.
-propose a logistic principal component analysis-based method to improve the representation of networks in [27].
-Meanwhile, evolutionary algorithms have also been applied to address the problem of community deception.
-Chen et al.
-use the genetic algorithm to minimize the modularity after the modification on edges in [28].
-Further, they categorize community deception based on the scales as community deception on node communities, target community and the entire community structure in [29].
-A self-adaptive evolutionary framework is then proposed to facilitate the search of the optimal edge set and a permanence-based method is developed to reduce the search space in [30].
-To alleviate the problem of the evaluation expense, Zhao and Cheong [31] propose a divide-and-conquer strategy to split the graph into subgraphs and optimize them with the genetic algorithm individually.
-Moreover, evolutionary algorithms have been applied to various other aspects of complex networks, including dynamic community detection [32], and module identification [33].
-These applications demonstrate the versatility and effectiveness of evolutionary algorithms in tackling diverse challenges within the field of complex network analysis.
-To date, considerable efforts have been directed towards tackling multi-task problems using evolutionary algorithms.
-In [5], Gupta et al.
-propose a paradigm, multi-factorial evolutionary algorithm, which can optimize different tasks simultaneously and has laid the foundation for the related research.
-For example, the optimization on single objective has been extended to the case of multiple objectives in [34], adopting the implicit transfer mode where the crossover is operated on the solutions aiming at different tasks [35, 36].
-There is another route referred to explicit transfer that deliberately incorporates task-specific information to guide knowledge exchange between tasks by using statistic information [37] or constructing a mapping matrix [38, 39].
-Beyond traditional fields, MTEO has been exploited within complex networks, such as in the optimization of network robustness.
-In [22], Wang et al.
-utilize the correlation of tasks to develop a new crossover operator, thereby obtaining knowledge from different individuals.
-As indicated in [40], structural destruction and cascading failure, two seemingly independent objectives, can promote each other in the optimization by developing a graph neural network-based method to exchange information between solutions aiming at different tasks.
-In one study of influence maximization [41], each task is treated as a transformation by formula, enabling the application of MTEO in this problem.
-In [21], Wu et al.
-promote the reconstruction of multiplex networks using MTEO with the inherent correlation of different layers.
-Similar work can be found in [42], in which the optimization of modularity in each layer is modeled as an individual task, thereby facilitating the community detection in multiplex networks.
-Despite that networks often share the same characteristics, cross-domain evolutionary optimization, especially concerning network structures, has received limited attention.
-Motivated by this overlooked potential, we introduce a novel optimization framework—Multi-Domain Evolutionary Optimization.
+[22] utilized the correlation of tasks to develop a new crossover operator, thereby injecting knowledge into different individuals.
+According to [47], a graph neural network-based method has been developed that facilitates the exchange of information, demonstrating that structural destruction and cascading failure, though seemingly unrelated goals, can mutually enhance optimization outcomes.
+In a study on influence maximization [48], each task is formulated as a transformation, allowing for the effective application of MTEO and resulting in enhanced performance.
+Wu et al.
+[21] promoted the reconstruction of multiplex networks using MTEO with the inherent correlation of different layers.
+Similar research by Lyu et al.
+[50] formulates the modularity optimization within each layer as an independent task, facilitating community detection in multiplex networks.
+Even though networks typically display common patterns, there has been minimal focus on cross-domain evolutionary optimization, especially with respect to network structures.
+Motivated by this overlooked potential, we introduce a novel optimization framework—multi-domain evolutionary optimization (MDEO).
+The comparison of the proposed MDEO and the existing MTEO literature is shown in Table I.
 III Optimization problem formulation
-Community deception involves nodes intentionally concealing their true community affiliations by modifying (adding and removing) connections.
-In this study, we select this challenging combinatorial problem [43, 44] in complex networks as an illustration of our MDEO for the following concerns:
+As a security-related task [51, 52, 53], community deception involves nodes intentionally concealing their true community affiliations by modifying (adding and removing) connections, which serves as the illustration of our MDEO for the following concerns:
 Exploiting Community Detection Algorithms: With the assistance of community detection algorithms, densely interconnected groups of nodes with high intra-group connectivity and low inter-group connectivity can be easily identified [50, 54].
-The development of community detection, while useful for understanding network structures and interactions, can also have certain dangerous aspects caused by the potential misuse or exploitation of community information, as suggested in [46, 47].
+The development of community detection, while useful for understanding network structures and interactions, can also have certain dangerous aspects caused by the potential misuse or exploitation of community information [55, 56].
 Complexity of Community Deception: As an edge-level task, the search space of community deception is $|V|^{2}$, much higher than node-level tasks of $|V|$.
@@
 Community deception serves as an innovative advancement in network science, radically altering node affiliations through minimal topological adjustments.
-Specifically, the modification on the network involves the addition of nonexistent edges and the deletion of existing edges.
-Let $G=\{V,E\}$ be a network where $V$ and $E$ denote the nodes and edges, respectively.
-The updated edge set can be defined as follows:
-$E^{\prime}=\left(E\cup E^{+}\right)\backslash E^{-},$ (1)
-where $E^{+}$ and $E^{-}$ refers to the addition and deletion of edges, that is:
+Specifically, the modification of the network involves the addition of nonexistent edges and the deletion of existing edges.
+Suppose $G=\{V,E\}$ is a network where $V$ and $E$ denote the nodes and edges, respectively.
+The modified edges can be defined as follows:
+[…]
+where $E^{+}$ and $E^{-}$ denote the edges to add and delete, i.e.,
 $\begin{array}[]{l}E^{+}\subseteq\{(u,v):u\in V\wedge v\in V,(u,v)\notin E\},\\ E^{-}\subseteq\{(u,v):u\in V\wedge v\in V,(u,v)\in E\},\end{array}$ (2)
@@
 $\left|E^{+}\right|+\left|E^{-}\right|\leq\beta,$ (3)
-where $\beta$ refers to the perturbance budget, meaning the number of edges that can be modified.
-Assume $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}^{\prime}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$ are the community partitions of original network $G$ and modified network $G^{\prime}$, then community deception can be represented as a combinatorial problem:
+where $\beta$ is the perturbance budget, referring to the maximum number of edges to rewire.
+Assume there are two community structures belonging to the original network $G$ and modified network $G^{\prime}$, $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}^{\prime}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, respectively, the problem of community deception is defined as:
 $\underset{\left\{E^{+},E^{-}\right\}}{\arg\max}\hskip 5.0pt\{\phi(\widetilde{\mathcal{C}},\widetilde{\mathcal{C}^{\prime})},E^{+},E^{-}\}.$ (4)
-The function $\phi$ refers to the difference between the community structures, before and after modification.
-IV Multi-Domain evolutionary optimization on community deception
-In this section, the proposed MDEO will be introduced in detail, and its diagram can be found in Figure 1.
+The function $\phi$ refers to the disparity between the community structures before and after modification. To validate the generalizability of MDEO, the node-level task is also tested in Section V-F.
+IV Multi-Domain evolutionary optimization
+This section provides a detailed introduction to the proposed MDEO, with its diagram shown in Figure 1. The mathematical notations used in MDEO are listed in Table II.
+TABLE II: Mathematical notations.
+  Symbols Definition
+$G$ A network consisting of nodes $V$ and edges $E$
+$\mathcal{G}$ Set of networks for optimization
+$\widetilde{\mathcal{C}}$ Community distribution of the network $G$
+$\beta$ Modification budget for the network $G$
+$\mathcal{A}$ Set of anchor nodes in the networks
+$v^{X,l}_{i}$ The $i$-th largest-degree node in the node set $X$
+$v^{X,s}_{i}$ The $i$-th smallest-degree node in the node set $X$
+$\Phi^{i\rightarrow j}$ Embedding mapping from network $G_{i}$ to network $G_{j}$
+$M_{i\rightarrow j}$ Edge mapping from network $G_{i}$ to network $G_{j}$
+$\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}$ Aligned community between networks $G_{i}$ and $G_{j}$
+$\mathcal{S}^{\mathcal{G}}_{i,j}$ Similarity bwtween networks $G_{i}$ and $G_{j}$
+$T_{i\leftarrow j}$ Solution transferred from network $G_{j}$ to network $G_{i}$
 Figure 1: The diagram of multi-domain evolutionary optimization.
-Four networks representing different systems are optimized simultaneously.
+Four networks representing different systems (brain network, power network, social network, communication network) are optimized simultaneously.
 The transfer of knowledge across networks is achieved through the trained network alignment model.
-$e^{i}_{j}$ denote the $j$th edge in $i$th network and $M_{i,j}$ is the edge mapping from $i$th network to $j$th network.
-The solution with the colors of yellow and light red refers to the elite solution and transferred solution from other networks, respectively.
+$e^{i}_{j}$ denote the $j$-th edge in $i$-th network and $M_{i\rightarrow j}$ is the edge mapping from the $i$-th network to the $j$-th network.
+The solution shown in yellow represents the transferred solution, while the light red color represents the elite solution from other networks.
 The red edge is the mutated edge that will be replaced by the edge of elite solutions of other networks.
+The knowledge exchange occurs when a series of conditions are met, and only elite solutions will be utilized to assist the optimization of other networks.
 IV-A Graph similarity
-The measurement of graph similarity involves the consideration of the interactions between nodes and higher-order structures, enabling us to perform selective knowledge transfer between networks.
-Instead of transferring all available information, it allows us to focus exclusively on pertinent domains, minimizing the risk of negative transfer and reducing the computational complexity.
+The measurement of graph similarity enables us to perform selective knowledge transfer between networks.
+Instead of transferring all available knowledge, it allows us to focus exclusively on pertinent domains, minimizing the risk of negative transfer and reducing computational complexity.
 In general, the graph similarity can be easily obtained by averaging node similarity of networks [57] or computing the distance of graph embeddings [58, 59].
 However, the former approach oversimplifies the complex interactions within the graph, ignoring the community structure and global patterns of networks.
-On the other hand, existing graph pooling techniques are typically trained on small graphs, potentially constraining their efficacy in accurately representing large-scale graphs.
-In addition to overcoming the above problems, this similarity is also expected to facilitate the construction of training data for the network alignment model.
-To this end, we propose a community-based measurement of graph similarity.
+On the other hand, existing graph pooling techniques are typically trained on small graphs, potentially constraining their efficacy on large-scale graphs.
+Here, we propose a community-level measurement of graph similarity.
 Given a network $G=(V,E)$ and its community distribution $\widetilde{\mathcal{C}}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\right\}$, for each community $\mathcal{C}\in\widetilde{\mathcal{C}}$, we can define $D_{\mathcal{C}}=\{d_{v}|v\in\mathcal{C}\}$ as the degree distribution of community $\mathcal{C}$ where $d_{v}$ is the node degree of $v$.
@@
 In addition, another challenge emerges due to the varying lengths of these normalized distributions across networks, which complicates the direct comparison of community degree distributions.
-To address this issue, we introduce a new definition that unifies the degree distributions across communities of different scales into the same length.
+To address this issue, we unify the degree distributions across communities of different scales into the same length.
 Definition 1.
 Degree interval-based distribution: Given a community $\mathcal{C}\in\widetilde{\mathcal{C}}$, then we can obtain the normalized degree distribution $D^{\prime}_{\mathcal{C}}=\left\{d^{\prime}_{1},d^{\prime}_{2},\ldots,d^{\prime}_{n}\right\}$ where $n$ is the number of nodes in community.
-Let $I_{k}=\left[a_{k},b_{k}\right)$ represent the $k$th interval, where $a_{k}$ is the lower bound, and $b_{k}$ is the upper bound of the interval, then we can have
+Let $I_{k}=\left[a_{k},b_{k}\right)$ represent the $k$-th interval, where $b_{k}$ is the upper bound of the interval and $a_{k}$ is the lower bound, then we can have
 $I_{j}=[(b_{j}-a_{j})*j,(b_{j}-a_{j})*(j+1)).$ (5)
 For $x_{i}\in D^{\prime}_{\mathcal{C}}$, corresponding interval $I_{j}$ can be found such that: $x_{i}\in I_{j}$ where $j=\left\lfloor\frac{x_{i}}{b_{j}-a_{j}}\right\rfloor$.
-Then, we can calculate the number of data points $x_{i}\in D^{\prime}_{\mathcal{C}}$ that fall into each interval $I_{k}$: $h_{j}=\operatorname{Count}\left(x_{i}\right.$ such that $\left.x_{i}\in I_{j}\right)$ for each $j$.
+Then, we can calculate the number of data points $x_{i}\in D^{\prime}_{\mathcal{C}}$ that fall into each interval $I_{k}:$ $h_{j}=\operatorname{Count}\left(x_{i}\right$.
+such that $\left.x_{i}\in I_{j}\right)$ for each $j$.
 Here, we set $b_{j}-a_{j}=0.2$ for $j\in\{0,1,2,3,4\}$.
@@
 $\operatorname{Diff}\left(\mathcal{C},\mathcal{C}^{\prime}\right)$ $=\frac{1}{2}\left(\mathbb{KL}\left(H_{\mathcal{C}}\|H_{\mathcal{C}^{\prime}}\right)+\mathbb{KL}\left(H_{\mathcal{C}^{\prime}}\|H_{\mathcal{C}}\right)\right)$
-×$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|),$ (7)
-where $\mathbb{KL}(.)$ refers to Kullback–Leibler divergence and the second item $\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|)$ is used to evaluate the difference of the scales of communities.
-Traditional graph representations focus on pairwise relationships between nodes while high-order graphs extend this concept and it is able to capture interactions involving more than two nodes.
-Therefore, we also consider the high-order information of networks, allowing for a more refined comparison that captures structural patterns.
+$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|),$ (7)
+where $\mathbb{KL}(.)$ refers to the Kullback–Leibler divergence and the second item $\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|)$ is used to quantify the disparity in community sizes.
+Traditional graph representations focus on pairwise relationships between nodes while higher-order graphs extend this concept and can capture interactions involving more than two nodes.
+Therefore, we also take into account higher-order network knowledge, enabling a more refined analysis of structural information.
 Given a network $G$ and a motif $M$, the network $G$ can be converted into a weighted network $G_{M}$ as follows:
 ${G}^{{M}}=\left\{{V},{E}^{{M}}\right\},$ (8)
-where ${E}^{M}\in E$ represents the set of edges within ${G}^{M}$.
-We can derive the adjacency matrix of hypergraph, denoted as $A_{M}$, in which each entry $(A_{M})_{ij}$ corresponds to the count of motifs that include the edge $e_{ij}$.
-We can measure the high-order difference between $\mathcal{C}$ and $\mathcal{C}^{\prime}$ as follows:
+where ${E}^{M}\in E$ represents the set of edges of ${G}^{M}$. We can measure the higher-order difference between $\mathcal{C}$ and $\mathcal{C}^{\prime}$ as follows:
 $\operatorname{Diff}^{M}\left(\mathcal{C},\mathcal{C}^{\prime}\right)$ $=\frac{1}{2}\left(\mathbb{KL}\left(H^{M}_{\mathcal{C}}\|H^{M}_{\mathcal{C}^{\prime}}\right)+\mathbb{KL}\left(H^{M}_{\mathcal{C}^{\prime}}\|H^{M}_{\mathcal{C}}\right)\right)$
-×$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|).$ (9)
-The similarity between communities is then defined as follows:
-$\mathcal{S}(\mathcal{C},\mathcal{C}^{\prime})=\textbf{Exp}(-\operatorname{Diff}\left(\mathcal{C},\mathcal{C}^{\prime}\right))+\textbf{Exp}(-\operatorname{Diff}^{M}\left(\mathcal{C},\mathcal{C}^{\prime}\right))$ (10)
+$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|).$ (9)
+The similarity between communities is defined as follows:
+$\mathcal{S}(\mathcal{C},\mathcal{C}^{\prime})=\textbf{Exp}(-\operatorname{Diff}\left(\mathcal{C},\mathcal{C}^{\prime}\right))+\textbf{Exp}(-\operatorname{Diff}^{M}\left(\mathcal{C},\mathcal{C}^{\prime}\right)).$ (10)
 Definition 2.
-Aligned community: Given two networks $G_{A}$, $G_{B}$ and their corresponding community structures $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, we can have the community-wise similarity matrix $\mathcal{S}\in\mathbb{R}^{k\times k^{\prime}}$ where $\mathcal{S}_{i,j}=\mathcal{S}(\mathcal{C}_{i},\mathcal{C}^{\prime}_{j}),i\in\{1,2,\cdots,k\},j\in\{1,2,\cdots,k^{\prime}\}$ Then, we define the set of aligned communities as follows:
-$\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}=\{\{\mathcal{C}_{a_{1}},\mathcal{C}^{\prime}_{b_{1}}\},\{\mathcal{C}_{a_{2}},\mathcal{C}^{\prime}_{b_{2}}\},\cdots,\{\mathcal{C}_{a_{z}},\mathcal{C}^{\prime}_{b_{z}}\}\},$ (11)
-where $z=\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)$.
-$\mathcal{C}_{a_{i}}$ and $\mathcal{C}^{\prime}_{b_{i}}$ refer to the $a_{i}$th and $b_{i}$th community in $\widetilde{\mathcal{C}_{A}}$ and $\widetilde{\mathcal{C}_{B}}$, respectively.
-Accordingly, we can have the set of their corresponding similarity values as follows:
-$\mathcal{S}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}=\{\mathcal{S}_{a_{1},b_{1}},\mathcal{S}_{a_{2},b_{2}},\cdots,\mathcal{S}_{a_{z},b_{z}}\},$ (12)
+Aligned community: Given two networks $G_{A}$, $G_{B}$ and their corresponding community structures $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, we can have the community-wise similarity matrix $\mathcal{S}_{i,j}\in\mathbb{R}^{k\times k^{\prime}}$ where $i\in\{1,2,\cdots,k\},j\in\{1,2,\cdots,k^{\prime}\}$.
+Then, we define the set of aligned communities as follows:
+$\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}=\{\{\mathcal{C}_{a_{1}},\mathcal{C}^{\prime}_{b_{1}}\},\{\mathcal{C}_{a_{2}},\mathcal{C}^{\prime}_{b_{2}}\},\cdots,\{\mathcal{C}_{a_{z}},\mathcal{C}^{\prime}_{b_{z}}\}\},$ (11)
+[…]
+$\mathcal{S}^{\widetilde{\mathcal{C}}}_{A,B}=\{\mathcal{S}_{a_{1},b_{1}},\mathcal{S}_{a_{2},b_{2}},\cdots,\mathcal{S}_{a_{z},b_{z}}\},$ (12)
 which satisfies
@@
 The similarity between $G_{A}$ and $G_{B}$ is defined as the average similarity of aligned community pairs.
-$\mathcal{S}^{\mathcal{G}}_{A,B}=\frac{1}{\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)}\sum_{x\in\mathcal{S}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}}x.$ (14)
-The process of calculating the graph similarity can be found in Algorithm 1.
+$\mathcal{S}^{\mathcal{G}}_{A,B}=\frac{1}{\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)}\sum_{x\in\mathcal{S}^{\widetilde{\mathcal{C}}}_{A,B}}x.$ (14)
+The graph similarity matrix $\mathcal{S}^{\mathcal{G}}$ is initialized based on Eq.
+14.
+Initially, it is used to identify the assisted network for each network and is subsequently updated to manage the transfer of solutions from each assisted network.
+The process of calculating the graph similarity is presented in Algorithm 1.
 Algorithm 1 Graph Similarity
 Input: Networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$
-Output: The quantitative similarity $S_{A,B}$ of $G_{A}$ and $G_{B}$.
-1:Obtain high-order graphs ${G}^{m}_{A}=({V}_{A},{E}^{m}_{A})$ and ${G}^{m}_{B}=(V_{B},{E}^{{m}}_{B})$
+Output: The graph similarity between $G_{A}$ and $G_{B}$
+1:Obtain higher-order graphs ${G}^{M}_{A}=({V}_{A},{E}^{M}_{A})$ and ${G}^{M}_{B}=(V_{B},{E}^{{M}}_{B})$
 2:Identify the community of $G_{A}$ and $G_{B}$ to obtain $\widetilde{\mathcal{C}}_{A}=\{\mathcal{C}_{1},\mathcal{C}_{2}\,\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}}_{B}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2}\,\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$
@@
 6:for $i=1$ to $\mathbf{Min}\{k,k^{\prime}\}$ do
-7:     Find the largest value of $\mathcal{S}$ and its two corresponding communities as well as its index $i_{\text{min}},j_{\text{min}}=\text{argmin}S_{i,j}$ where $1\leq i\leq k,1\leq j\leq k^{\prime}$
+7:  Find the largest value of $\mathcal{S}$ and its two corresponding communities as well as its index $i_{\text{max}},j_{\text{max}}=\text{argmax}S_{i,j}$ where $1\leq i\leq k,1\leq j\leq k^{\prime}$
 8:  Update $\mathcal{S}^{\mathcal{G}}_{A,B}\leftarrow\mathcal{S}^{\mathcal{G}}_{A,B}+\mathbf{Max}(\mathcal{S})$
-9:     Make the $i_{\text{min}}$ row and $j_{\text{min}}$ column of $S$ as 00
+9:  Set the $i_{\text{max}}$ row and $j_{\text{max}}$ column of $S$ to 0
 10:end for
-11:Normalize the similarity $\mathcal{S}^{\mathcal{G}}_{A,B}$ based on the number of communities of $G_{A}$ and $G_{B}$
+11:Normalize the similarity $\mathcal{S}^{\mathcal{G}}_{A,B}$
 IV-B Graph embedding
-To construct the node mappings across different networks, we employ graph autoencoders (GAE) to generate embeddings since it can capture meaningful and compact graph representations without the labeled data [51].
+To construct the node mappings across different networks, we employ graph autoencoders (GAE) to generate embeddings capturing meaningful and compact graph representations without the labeled data [60].
 In GAE, the graph convolutional network (GCN) is taken as the encoder [61], i.e.,
 $\mathbf{Z}=f_{encoder}(G)=\operatorname{GCN}(\mathbf{X},\mathbf{A}),$ (15)
-where $\mathbf{A}$ denotes the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{N\times d}$ denotes nodes features. The information aggregation of GCN is defined as follows:
+where $\mathbf{A}$ denotes the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{N\times d}$ denotes nodes features. The information aggregation of GCN is
 $\operatorname{GCN}(\mathbf{X},\mathbf{A})=\widetilde{\mathbf{A}}\operatorname{ReLU}\left(\widetilde{\mathbf{A}}\mathbf{X}\mathbf{W}_{\mathbf{0}}\right)\mathbf{W}_{\mathbf{1}},$ (16)
 where $\mathbf{W}_{\mathbf{0}}$ and $\mathbf{W}_{\mathbf{1}}$ are learnable parameters and $\widetilde{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$.
-The decoder takes the latent space representation $\mathbf{Z}$ as the input to reconstruct the adjacency matrix. The inner-product is adopted as the decoder function:
+The decoder takes the latent space representation $\mathbf{Z}$ as input to reconstruct the adjacency matrix. The inner product is adopted as the decoder function:
 $\hat{\mathbf{A}}=f_{decoder}(\mathbf{Z})=\sigma\left(\mathbf{Z}\mathbf{Z}^{\mathrm{T}}\right),$ (17)
@@
 GAE is trained by minimizing the reconstruction loss that measures the dissimilarity between the adjacency matrix $\mathbf{A}$ and the reconstructed matrix $\hat{\mathbf{A}}$. The loss function is:
-$\mathcal{L}_{reconstruct}=-\frac{1}{N}\sum y\log\hat{y}+(1-y)\log(1-\hat{y}),$ (18)
+[…]
 where $y$ refers to the element in adjacency matrix $\mathbf{A}$ and $\hat{y}$ is the element that has the same position with $y$ in $\hat{\mathbf{A}}$.
@@
 We also denote the inverse mapping in a similar manner as $\Phi^{b\rightarrow a}\left(\mathbb{E}_{B}\left(v\right)\right)=\mathbb{E}_{A}\left(u\right)$.
-In the supervised learning component, a set of anchor nodes is carefully selected in terms of the structural role in both networks, serving as labels to train the network alignment model.
-Given two networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$, we can then obtain their aligned community pairs $\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}$ as in Definition 2, which is the basis of the construction of anchor nodes.
-To make the network alignment model adapted to the cross-domain scenario, we consider both the important nodes and trivial nodes in each aligned community pair as anchor nodes.
+A collection of anchor nodes is carefully chosen based on their structural functions in both networks, serving as labels to train the network alignment model.
+Given two networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$, we can then obtain their aligned community pairs $\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}$ as in Definition 2, which is the basis of the construction of anchor nodes.
+To adapt the network alignment model for multi-domain scenarios, both trivial and important nodes within each aligned community pair will be considered as anchor nodes.
 Definition 4.
-Anchor node (large degree) – $\mathcal{A}^{large}$: Let $v^{X,l}_{i}$ denote the $i$th largest-degree node in node set $X$.
-Then, given two networks $G_{A}$ and $G_{B}$, and an aligned community pair $\{\mathcal{C},\mathcal{C}^{\prime}\}\in\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}$, the set of anchor node of large degree is denoted as follows:
-$\mathcal{A}^{large}=\bigcup\limits^{\{\mathcal{C},\mathcal{C}^{\prime}\}\in\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}\{v^{\mathcal{C},l}_{i},v^{\mathcal{C}^{\prime},l}_{i}\},i\in\{1,...,k_{l}\},$ (19)
+Anchor node (large degree) – $\mathcal{A}^{large}$: Suppose the $i$-th largest-degree node in the set $X$ is $v^{X,l}_{i}$.
+Then, given two networks $G_{A}$ and $G_{B}$, and an aligned community pair $\{\mathcal{C},\mathcal{C}^{\prime}\}\in\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}$, the set of anchor large-degree nodes can be formulated as follows:
+$\mathcal{A}^{large}=\bigcup\limits^{\{\mathcal{C},\mathcal{C}^{\prime}\}\in\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}}\{v^{\mathcal{C},l}_{i},v^{\mathcal{C}^{\prime},l}_{i}\},i\in\{1,...,k_{l}\},$ (19)
 where $k_{l}=\mathbf{Min}\{log_{2}|\mathcal{C}|,log_{2}|\mathcal{C}^{\prime}|\}$.
 The loss for aligning the important nodes can be obtained as follows:
-$\begin{split}\mathcal{L}_{large}&=\sum\limits_{\{u,v\}\in\mathcal{A}^{large}}dist(\Phi^{a\rightarrow b}(\mathbb{E}(u)),\mathbb{E}(v))\\ &\quad+dist(\Phi^{b\rightarrow a}(\mathbb{E}(v)),\mathbb{E}(u)).\end{split}$ (20)
-where $dist$ refers to mean square error (MSE), measuring the difference between two embeddings.
+$\begin{split}\mathcal{L}_{large}&=\sum\limits_{\{u,v\}\in\mathcal{A}^{large}}dist(\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u)),\mathbb{E}_{B}(v))\\ &\quad+dist(\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v)),\mathbb{E}_{A}(u)),\end{split}$ (20)
+where distdist refers to the mean square error (MSE), measuring the difference between two embeddings.
 To preserve the structural information, we will not consider the nodes with the least degree in the community.
@@
 Definition 5.
-Anchor node (small degree) – $\mathcal{A}^{small}$: Let $v^{X,s}_{i}$ denote the $i$th least-degree node in the node set $X$.
-Given an aligned pair of nodes $\{u,v\}\in\mathcal{A}^{large}$, we can define
+Anchor node (small degree) – $\mathcal{A}^{small}$: Suppose the $i$-th least-degree node in the set $X$ is $v^{X,s}_{i}$.
+Given a set of aligned large-degree nodes $\{u,v\}\in\mathcal{A}^{large}$, the aligned small-degree nodes can be defined as follows:
 $\mathcal{A}^{small}=\bigcup\limits^{\{u,v\}\in\mathcal{A}^{large}}\{v^{\mathcal{N}(u),s}_{i},v^{\mathcal{N}(v),s}_{j}\},i,j\in\{1,...,k_{s}\},$ (21)
-where $k_{s}=\mathbf{Min}\{log_{2}|\mathcal{N}(u)|,log_{2}|\mathcal{N}(v)|\}$ and $\mathcal{N}(.)$ refers to the neighbor nodes.
+where $k_{s}=\mathbf{Min}\{log_{2}|\mathcal{N}(u)|,log_{2}|\mathcal{N}(v)|\}$ and $\mathcal{N}(.)$ denotes the neighbor nodes.
 These least important nodes are trivial in the network (and have little influence on the network structure), therefore, their roles are basically the same and will not require a one-to-one relationship.
@@
 Given two networks of $G_{A}$ and $G_{B}$, the loss function for training the mappings $\Phi^{a\rightarrow b}$ and $\Phi^{b\rightarrow a}$ is obtained as follows:
-$\mathcal{L}_{alignment}\left(\mathbf{W}_{s,t},\mathbf{b}_{s,t},\mathbf{W}_{t,s},\mathbf{b}_{t,s}\right)=\mathcal{L}_{large}+\mathcal{L}_{small}+\mathcal{L}_{us}.$ (24)
-Upon the optimization on $\mathcal{L}_{alignment}$, the trained mapping can align the remaining nodes between the two networks, enabling the solutions between networks to be transferred to each other.
+$\mathcal{L}_{alignment}\left(\mathbf{W}_{a,b},\mathbf{b}_{a,b},\mathbf{W}_{b,a},\mathbf{b}_{b,a}\right)=\mathcal{L}_{large}+\mathcal{L}_{small}+\mathcal{L}_{us}.$ (24)
+After optimizing $\mathcal{L}_{alignment}$, the trained mapping will align the remaining nodes between the two networks, facilitating the transfer of solutions between them.
+To establish node-level alignment between two networks based on their learned embeddings, we first project the embeddings from one network into the latent space of the other using learned mapping functions.
+Specifically, for each node $u\in V_{A}$, its embedding is transformed into the space of $G_{B}$ using the mapping function $\Phi^{A\rightarrow B}$.
+The node $v\in V_{B}$ whose embedding is closest, in terms of Euclidean distance, to the mapped embedding of node $u$ is identified, and its index is stored in the alignment mapping as $M_{A\rightarrow B}(u)=v$.
 Algorithm 2 Network Alignment
 Input: Networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$
-Output: Graph Mapping model $\mathbb{M}_{AB}$ and $\mathbb{M}_{BA}$
+Output: Graph Mappings $\Phi^{a\rightarrow b}$ and $\Phi^{b\rightarrow a}$
 1:Obtain the embeddings of $V_{A}$ and $V_{B}$ as $\mathbb{E}_{A}=\mathbf{GAE}(G_{A})$ and $\mathbb{E}_{B}=\mathbf{GAE}(G_{B})$
@@
 3:  #Unsupervised learning where $u\in V_{A}$ and $v\in V_{B}$
-4:     Map the embedding $\mathbb{E}_{A}(u)$ to the space of $G_{B}$ via $\mathbb{M}_{AB}$, i.e., $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then map the embeddings of $G_{B}$ back to the space of $G_{A}$ via $\mathbb{M}_{BA}$
+4:  Map the embedding $\mathbb{E}_{A}(u)$ to the space of $G_{B}$ via $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then map the embeddings of $G_{B}$ back to the space of $G_{A}$
 5:  Calculate the difference between $\Phi^{b\rightarrow a}(\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$ and $\mathbb{E}_{A}(u)$
-6:     Map the embedding $\mathbb{E}_{B}(v)$ to the space of $G_{A}$ via $\mathbb{M}_{BA}$, i.e., $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$, and then map the embeddings of $G_{A}$ back to the space of $G_{B}$ via $\mathbb{M}_{AB}$
+6:  Map the embedding $\mathbb{E}_{B}(v)$ to the space of $G_{A}$ via $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$, and then map the embeddings of $G_{A}$ back to the space of $G_{B}$
 7:  Calculate the difference between $\Phi^{a\rightarrow b}(\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$ and $\mathbb{E}_{B}(v)$
-8:     Calculate the unsupervised loss between the original embeddings and the automapped embeddings.
+8:  Calculate the unsupervised loss between the original embeddings and the automapped embeddings
 9:  #Supervised learning where $u\in V_{A}$ and $v\in V_{B}$, and $(u,v)$ is an anchor pair
-10:     Map $u$ to the space of $G_{B}$ via $\mathbb{M}_{AB}$ to obtain $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then calculating its difference with $\mathbb{E}_{A}(u)$
-11:     Map $v$ to the space of $G_{A}$ via $\mathbb{M}_{BA}$ to obtain $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$ and then calculating its difference with $\mathbb{E}_{B}(v)$
+10:  Map $u$ to the space of $G_{B}$ to obtain $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and calculating its difference with $\mathbb{E}_{B}(v)$
+11:  Map $v$ to the space of $G_{A}$ to obtain $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$, and calculating its difference with $\mathbb{E}_{A}(u)$
 12:  Use the anchor nodes to calculate the supervised loss
 13:end for
-14:Update the parameters of mappings $\mathbb{M}_{AB}$ and $\mathbb{M}_{BA}$
+14:Update the parameters of mappings $\Phi^{a\rightarrow b}$ and $\Phi^{b\rightarrow a}$
 IV-D Many-network evolutionary optimization
-In the evolutionary optimization of community deception, the chromosome is represented by the combination of genes (edges) with a parameter $\rho$ differentiating the edges to add or delete.
-The population of network $G_{i}\in\mathcal{G}$ can be defined as follows:
-$P_{i}=\left[\begin{array}[]{cccccc}e_{1,1}^{i}&e_{1,2}^{i}&\cdots&\cdots&e_{1,\beta_{i}}^{i}&\rho_{1}^{i}\\ e_{2,1}^{i}&e_{2,2}^{i}&\cdots&\cdots&e_{2,\beta_{i}}^{i}&\rho_{2}^{i}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ e_{n,1}^{i}&e_{n,2}^{i}&\cdots&\cdots&e_{n,\beta_{i}}^{i}&\rho_{b}^{i}\end{array}\right],$ (25)
+In evolutionary optimization for community deception, the chromosome is represented by a combination of genes (edges). The population of network $G_{i}\in\mathcal{G}$ can be defined as follows:
+$P_{i}=\left[\begin{array}[]{cccccc}e_{1,1}^{i}&e_{1,2}^{i}&\cdots&\cdots&e_{1,\beta_{i}}^{i}&\rho_{1}^{i}\\ e_{2,1}^{i}&e_{2,2}^{i}&\cdots&\cdots&e_{2,\beta_{i}}^{i}&\rho_{2}^{i}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ e_{n,1}^{i}&e_{n,2}^{i}&\cdots&\cdots&e_{n,\beta_{i}}^{i}&\rho_{n}^{i}\end{array}\right],$ (25)
 where the first $\rho$ edges are to add, and the rest $\beta-\rho$ are the edges to delete.
@@
 The entry $m_{ij}$ in the matrix denotes the number of common elements between $\mathcal{C}_{i}$ and $\mathcal{C}^{\prime}_{j}$.
-To facilitate the disengagement of nodes from their initial community affiliations, we can formulate
-$F_{1}=-\sum_{i=1}^{|\widetilde{\mathcal{C}}|}\sum_{j=1}^{|\widetilde{\mathcal{C}}^{\prime}|}\frac{|\mathcal{C}_{i}|}{|V|}\left(\frac{m_{ij}}{|\mathcal{C}_{i}|}\log_{2}\frac{m_{ij}}{|\mathcal{C}_{i}|}\right).$ (26)
-The objective of Eq.
-26 is to foster decentralization within the community structure, albeit it may inadvertently result in numerous tiny communities.
-Ideally, the nodes from one community should integrate seamlessly into other communities without significantly altering the community sizes.
-To mitigate this extreme scenario, we introduce a robust alternative function that recalibrates the confusion matrix as follows:
-$m^{\prime}_{ij}=\frac{m_{ij}}{|\mathcal{C}^{{}^{\prime}}_{j}|}*\log_{2}|\mathcal{C}^{{}^{\prime}}_{j}|.$ (27)
-The first item, ${m_{ij}}/{|\mathcal{C}^{{}^{\prime}}_{j}|}$ refers to the proportion of the nodes of the original community $\mathcal{C}_{i}$ accounting for the new community $\mathcal{C}^{{}^{\prime}}_{j}$.
-The second term is utilized to prevent situations where a single node forms a community independently.
-Instead of resorting to computing a global indicator like entropy, we opt for an elegant yet effective approach by constraining its maximum value as follows:
-$F_{2}=e^{-max(m^{\prime}_{ij})}.$ (28)
-Then, the fitness function $F$ can be defined as follows:
-$F=F_{1}*F_{2}.$ (29)
-[…]
+To facilitate the disengagement of nodes from their initial community affiliations, we can formulate the fitness function as follows:
+$F=-\left(\sum_{i=1}^{|\widetilde{\mathcal{C}}|}\sum_{j=1}^{|\widetilde{\mathcal{C}}^{\prime}|}\frac{|\mathcal{C}_{i}|}{|V|}\left(\frac{m_{ij}}{|\mathcal{C}_{i}|}\log_{2}\frac{m_{ij}}{|\mathcal{C}_{i}|}\right)\right)*e^{-\max(m^{\prime}_{ij})},$ (26)
+where $m_{ij}$ denotes the number of common nodes between $\mathcal{C}_{i}$ and $\mathcal{C}^{\prime}_{j}$, and $m^{\prime}_{ij}=\frac{m_{ij}}{|\mathcal{C}^{{}^{\prime}}_{j}|}*\log_{2}|\mathcal{C}^{{}^{\prime}}_{j}|$.
+For a network $G\in\mathcal{G}$, considering the knowledge transferred from all the rest networks $\mathcal{G}/G$ is undesirable because 1) the optimization efficiency will be compromised with the increase in the networks participating in transfer; 2) the solution from networks that have a low similarity with the target network may bring negative transfer.
+Therefore, only the solutions from those networks similar to the target network will be transferred to the target population.
+Let $\mathcal{G}^{\mathcal{S}}_{i}$ be the set of top $\sqrt{|\mathcal{G}|}$ similar networks to network $G_{i}$, which satisfies
 $\begin{cases}|\mathcal{G}^{\mathcal{S}}_{i}|=\sqrt{|\mathcal{G}|},\\ \forall G_{j}\in\mathcal{G}^{\mathcal{S}}_{i},\forall G_{k}\notin\mathcal{G}^{\mathcal{S}}_{i}:\mathcal{S}^{\mathcal{G}}_{i,j}>\mathcal{S}^{\mathcal{G}}_{i,k}.\end{cases}$} (27)
-The transfer solution requires several rounds of reproduction to be integrated into the population of target domain.
+The transferred solution requires several rounds of reproduction to be integrated into the population of the target domain.
 Combined with the interest of efficiency, we thus take $k=5$ empirically as the interval to determine whether to transfer or not.
-Let the best fitness values on $G_{i}$ in the current generation and $k$ generations ago be $F^{i}_{n\_0}$ and $F^{i}_{n\_k}$, respectively.
+Let the best fitness values on $G_{i}$ in the current generation and $k$ generations ago be $F^{i}_{n_{0}}$ and $F^{i}_{n_{k}}$, respectively.
 Then, we can obtain the fitness difference between the current generation and $k$ generations ago and the fitness difference between $k$ and $2k$ generations ago as follows:
-$d_{1}^{i}=\left|F_{n\_0}^{i}-F_{n\_k}^{i}\right|,\quad d_{2}^{i}=\left|F_{n\_k}^{i}-F_{n\_2k}^{i}\right|.$ (31)
+$d_{1}^{i}=\left|F_{n_{0}}^{i}-F_{n_{k}}^{i}\right|,\quad d_{2}^{i}=\left|F_{n_{k}}^{i}-F_{n_{2k}}^{i}\right|.$ (28)
 Then, we define the transfer condition as follows:
-$\begin{cases}d_{1,n}^{i}<d_{2,n}^{i},\\ n\geq 2k,\\ n\bmod k=0.\end{cases}$ (32)
-The motivation of Eq.32 is that if the evolution speed is slower than the previous, then excellent solutions of other similar domains will be transferred in to facilitate the optimization.
+$\begin{cases}d_{1}^{i}<d_{2}^{i},\\ n\geq 2k,\\ n\bmod k=0.\end{cases}$ (29)
+The motivation of Eq. 29 is that if the evolution speed is slower than the previous, then excellent solutions of other similar domains will be transferred to facilitate the optimization.
 To better use the transfer budget, we propose a self-adaptive mechanism to determine the number of solutions to transfer from $G\in\mathcal{G}^{\mathcal{S}}_{i}$ to $G_{i}$.
-Each transfer will be evaluated by its contribution to the new population.
-We will observe the difference between the current elite population $P^{Elite}_{i,n\_0}$ and the elite population before the last transfer $P^{Elite}_{i,n\_l}$, and its intersection with the last transfer ($P^{Transfer}_{i\leftarrow j,n\_l}$) from $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$.
+Each transfer will be evaluated by its contribution to the new population by observing the difference between the current elite population $P^{Elite}_{i,n_{0}}$ and the elite population before the last transfer $P^{Elite}_{i,n_{l}}$, and its intersection with the last transfer ($T_{i\leftarrow j,n_{l}}$) from $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$.
 That is,
-$I_{i\leftarrow j}^{\mathcal{G}}=\frac{|(P^{Elite}_{i,n\_0}\setminus P^{Elite}_{i,n\_l})\bigcap P^{Transfer}_{i\leftarrow j,n\_l}|}{P^{Elite}_{i,n\_0}}.$ (33)
+$I_{i\leftarrow j}^{\mathcal{G}}=\frac{|(P^{Elite}_{i,n_{0}}\setminus P^{Elite}_{i,n_{l}})\bigcap T_{i\leftarrow j,n_{l}}|}{|P^{Elite}_{i,n_{0}}|},$ (30)
+where $T_{i\leftarrow j}=\{M_{i\leftarrow j}(s)\mid s\in S_{j}\},$ and $S_{j}$ is a set of selected solutions from network $G_{j}$, and $M_{i\leftarrow j}$ is the function mapping each solution from the domain of $G_{j}$ to that of $G_{i}$.
 The similarity matrix is then updated as follows:
 $\mathcal{S}^{\mathcal{G}}_{i,j}\leftarrow I_{i\leftarrow j}^{\mathcal{G}}+\mathcal{S}^{\mathcal{G}}_{i,j},$ (31)
-where $\mathcal{S}^{\mathcal{G}}_{i,j}$ is the similarity between $G_{i}$ and $G_{j}$.
-In this manner, we are able to control the ratio of transferred solutions from different networks, given a fixed budget of transfer $\beta^{Transfer}$.
+where $\mathcal{S}^{\mathcal{G}}_{i,j}$ refers to the similarity between $G_{i}$ and $G_{j}$, which can be used to control the ratio of transferred solutions from different networks, given a fixed number of transferred solutions $|T_{i}|$.
 The networks that are more similar to the target network should transfer more solutions, i.e.,
-$|P^{Transfer}_{i\leftarrow j}|=\beta^{Transfer}*\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (35)
-To better use the knowledge of other domains, we design a new mutation operator, which utilizes the elite edge of other networks to replace the edge in its own domain.
-Given a network $G_{i}\in\mathcal{G}$, the probability of the network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$ being selected to transfer is defined as follows:
-$|P^{M}_{i\leftarrow j}|=\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (36)
-Note that as different networks have different budgets $\beta$, the repairs to transferred solutions are necessary to ensure compatibility when the sizes of the transferred solution and the solution of the target network are not matched.
+$|T_{i\leftarrow j}|=|T_{i}|*\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (32)
+Elements in the similarity matrix corresponding to non-similar networks are set to zero ($\mathcal{S}^{\mathcal{G}}_{i,j}=0,G_{j}\notin\mathcal{G}^{\mathcal{S}}_{i}$) prior to optimization, whereas elements representing network similarity are normalized during each update.
+Let the population of network $G_{i}$ be $P_{i}$ and knowledge transfer occurs from similar networks $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$, and the transferred high-quality solution $T_{i\leftarrow j}$ from a network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$ is incorporated into the population to assist the optimization process.
+Therefore, the updated population is given by:
+$P_{i}=\begin{cases}P_{i}\cup\bigcup_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}T_{i\leftarrow j},&\text{if knowledge transfer occurs,}\\ P_{i},&\text{otherwise.}\end{cases}$} (33)
+To better utilize the knowledge of other domains, we introduce a new mutation mechanism that leverages elite edges from other networks to replace edges in the current domain.
+The probability of selecting a network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$ for knowledge transfer is defined as follows:
+$\mathbb{P}^{M}_{i\leftarrow j}=\frac{\mathcal{S}^{\mathcal{G}}_{i,j}}{\sum\limits_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}\mathcal{S}^{\mathcal{G}}_{i,j}}.$ (34)
+Given a network $\mathcal{G}_{i}$, the deletion and addition candidates are denoted by $E_{i}^{M,D}$ and $E_{i}^{M,A}$, respectively.
+When knowledge transfer is applied, these candidates are derived from the transferred solutions of assisted networks $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$, which are organized into addition ($T_{i\leftarrow j}^{A}$) and deletion ($T_{i\leftarrow j}^{D}$) sets.
+The mutation candidate sets are then updated accordingly as follows:
+$E_{i}^{M,A}=\bigcup_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}T_{i\leftarrow j}^{D},\quad E_{i}^{M,D}=\bigcup_{G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}}T_{i\leftarrow j}^{A},$} (35)
+where $T_{i\leftarrow j}^{A}$ represents the set of edges for addition, and $T_{i\leftarrow j}^{D}$ represents the set of edges for deletion from the transferred solution of network $G_{j}\in\mathcal{G}^{\mathcal{S}}_{i}$.
+As for the selection in our work, each individual has an equal probability of being selected, promoting diversity and helping prevent premature convergence on suboptimal solutions.
+Additionally, we adopt an elitist mechanism to ensure that the highest-quality solutions are preserved.
+As different networks have different budgets $\beta$, the repairs to transferred solutions are necessary to ensure compatibility when the sizes of the transferred solution and the solution of the target network are not matched.
 Given an assisted network $G_{A}$ and a target network $G_{T}$, if $\beta_{A}>\beta_{T}$, then the edges in the transferred solution will be removed randomly until the length of the transferred solution meets the requirement.
 Otherwise, the edges will be sampled from existing edges as deletion and nonexistent edges as addition to make the transferred solution valid.
-In the process of initialization, the chromosomes, denoting edge sets, undergo initialization through a sampling process, where both existing and nonexistent edges are included, with a random ratio.
-In selection, the roulette wheel strategy is utilized to select the chromosomes to participate in evolution.
-The chromosome with a larger fitness will be more likely to be selected.
-The elitism strategy is also adopted to safeguard against the interference of inferior genes, in which the top 10% of chromosomes with superior fitness are retained and directly propagated to the succeeding generation.
-Note that the operation of crossover and mutation may change the ratio of the added edge and deleted edge in the chromosome (edge set), thus the parameter $\rho$ will be updated during evolution.
+The pseudocode of the many-network evolutionary optimization can be found in Algorithm 3.
 Algorithm 3 Many-Network Evolutionary Optimization
 Input: Networks $G_{i}=(V_{i},E_{i})\in\mathcal{G}$
-Output: The optimized solution for the task on each network
+Output: The optimal solutions for the task across each network
 1:Initialize population for each network
-2:Initialize correlation coefficient matrix based on similarity matrix
+2:Initialize similarity matrix
 3:Identify the top-k similar networks for each network
@@
 11:     end for
-12:              Update and normalize correlation matrix
+12:     Update and normalize similarity matrix
 13:   end if
@@
 17:end while
+IV-E Complexity analysis
+[…]
+If the networks are optimized independently, the time complexity is $O(|\mathcal{G}|\cdot|P|\cdot N_{t})$, where $N_{t}$ denotes the number of evolutionary rounds, and $|P|$ is the population size in each network.
+In the MDEO framework, the solutions from $\log|\mathcal{G}|$ networks are transferred to assist others, but the number of solutions transferred, denoted by $|T|$, is fixed, and all assisted networks share this budget.
+Consequently, the worst-case complexity for MDEO could be $O(|\mathcal{G}|\cdot|P\cup T|\cdot N_{t})$.
+However, transfers only occur at least every five generations and additionally must satisfy the requirements of Eq.
+29.
+Therefore, the practical complexity will be lower than this worst-case scenario.
+Since $|T|$ is a constant and significantly smaller than $|P|$, the complexity of MDEO will only be marginally higher than that of separately optimized networks, i.e., single-domain evolutionary optimization (SDEO), thereby justifying the practical efficiency of MDEO.
 V Experimental studies
@@
 V-A Experimental setting
-We use eight different real-world networks [53] of various sizes from different domains to validate our proposed MDEO.
-In the process of evolution, the population size of each network is set to 100.
-In the reproduction, the probabilities of crossover and mutation are set to 0.5 and 0.1, respectively.
-In the comparison, the effectiveness of random modification will be investigated and the pioneering work in community deception, DICE [43] will also used as one of baselines.
-We also optimize the modularity that is the most recognized metric in the studies of community, using the genetic algorithm as in [28] to conceal community information.
-To demonstrate the effectiveness, the classical evolutionary optimization, referred to single-domain evolutionary optimization (SDEO), will also be included in the comparison.
-V-B Experiment on real-world networks
-In Table II(b), we compare the efficacy of different community deception algorithms by examining two metrics of NMI and ARI.
-The lower NMI and ARI are, the better protection the community deception algorithm achieves.
-We also use Wilcoxon rank sum tests to observe the statistical difference between the compared methods and MDEO, with the significance level $p=0.05$.
-The symbols ‘+’ and ‘-’ means the compared method is better/worse than MDEO and $\approx$’ means that the compared method has the similar performance with MDEO.
-The community deception algorithms are tested with three attackers (community detection algorithms) on eight networks of different domains.
-NMI and ARI basically exhibit the same trend but there are still some exceptions.
-For example, the NMI value of SDEO and MDEO is higher than DICE while their ARI values are much lower than that of DICE.
-As observed, our proposed MDEO clearly outperforms other methods in most cases and is better than SDEO, suggested by its lower average NMI and ARI than other compared methods.
+To obtain graph embeddings, we adopt a two-layer GCN as the encoder where the degree centrality, closeness centrality and community information are collected as node features.
+In the process of evolution, the population size of each network is set to 100, and the probabilities of crossover and mutation are set to 0.5 and 0.1, respectively.
+The number of transferred solutions is 30.
+In our study, we use two classical community detection algorithms, FastGreedy [62] and WalkTrap [63], as attackers to assess whether the community structure has been effectively obfuscated through community deception methods.
+The results are averaged over 20 independent runs and the implementation data is available online [^1] https://networkrepository.com/.
+V-B Benchmark
+The mainstream approaches to community deception include heuristic methods, GNN-based method, and evolutionary optimization techniques.
+Accordingly, some of each have been chosen as benchmarks, specifically: RAM: A heuristic approach that randomly rewires all possible edges in the network [51]; DICE: A heuristic approach that randomly deletes intra-community edges and adds inter-community edges [51]; Mod: This method uses the genetic algorithm to optimize modularity to modify the network edges to conceal community information [27]; CD-ATTACK: This method alters links by using constrained adversarial graphs generator based on graphs neural networks (GNNs) [64]; CoeCo: This approach employs a divide-and-conquer strategy for scalable evolutionary optimization.
+It optimizes both the subgraphs and the original graph cooperatively to identify the optimal edge set for community obfuscation [30]; SDEO: The classical evolutionary optimization without knowledge transfer, where each network is optimized separately, providing a baseline for comparison against our proposed MDEO.
+V-C Experiment on real-world networks
+To examine the efficacy of different community deception algorithms, we adopt two metrics of normalized mutual information (NMI) [65] and adjusted rand index (ARI) [66] to observe the change in community structure, shown in Table IV(b).
+The lower the NMI and ARI are, the better protection the community deception algorithm achieves.
+The statistical disparities between MDEO and other methodologies were assessed using the Wilcoxon rank sum test, setting the significance level at $p=0.05$.
+Symbols ‘+’ and ‘-’ are used to show whether the alternative methods outperform or underperform compared to MDEO, respectively, and $\approx$’ signals performance equivalence.
+The community deception algorithms are tested with two attackers (community detection algorithms) on eight networks of different domains.
+As observed, our proposed MDEO consistently outperforms other methods, including heuristic approaches (RANDOM, DICE), metaheuristic algorithms (MoD, CoeCo), and the GNN-based method (CD-ATTACK).
+It also demonstrates superior performance compared to SDEO, as indicated by its lower average NMI and ARI scores across the evaluated networks.
 This result indicates the effectiveness of our method in securing community information.
@@
 For example, the NMI and ARI in Adjnoun when dealing with the attack from FastGreedy, are lower than other networks.
-Network is also an important variable, as suggested by that NMI and ARI in Adjnoun are higher when dealing with WalkTrap than other networks.
-TABLE I: The comparison of the mean and standard deviation of NMI and ARI obtained from different community deception algorithms.
-The results are averaged from 20 independent runs.
-Three community detection algorithms are tested, i.e., FastGreedy and WalkTrap.
+Network is also an important variable, as suggested by that the NMI and ARI in Adjnoun are higher when dealing with WalkTrap than other networks.
+TABLE III: The comparison of the mean and standard deviation of NMI and ARI obtained from different community deception algorithms on eight tested networks.
+FastGreedy and WalkTrap are used to examine the difference in community structure before and after modification, respectively.
   Network NMI ARI
-RAM DICE Mod SDEO MDEO RAM DICE Mod SDEO MDEO
-Dolphins 0.75$\pm$0.060(-) 0.74$\pm$0.082(-) 0.76$\pm$0.094(-) 0.51$\pm$0.056(-) 0.43$\pm$0.041 0.70$\pm$0.072(-) 0.69$\pm$0.091(-) 0.71$\pm$0.115(-) 0.31$\pm$0.053(-) 0.24$\pm$0.032
-Lesmis 0.79$\pm$0.113(-) 0.75$\pm$0.118(-) 0.80$\pm$0.093(-) 0.55$\pm$0.024(-) 0.47$\pm$0.035 0.72$\pm$0.178(-) 0.65$\pm$0.187(-) 0.76$\pm$0.113(-) 0.31$\pm$0.026(-) 0.25$\pm$0.025
-Polbooks 0.85$\pm$0.100(-) 0.89$\pm$0.079(-) 0.92$\pm$0.060(-) 0.54$\pm$0.047($\approx$) 0.52$\pm$0.035 0.84$\pm$0.117(-) 0.89$\pm$0.093(-) 0.92$\pm$0.079(-) 0.45$\pm$0.072(-) 0.41$\pm$0.054
-Adjnoun 0.54$\pm$0.118(-) 0.42$\pm$0.094(-) 0.48$\pm$0.118(-) 0.24$\pm$0.036(-) 0.24$\pm$0.026 0.43$\pm$0.159(-) 0.29$\pm$0.119(-) 0.37$\pm$0.154(-) 0.09$\pm$0.018(-) 0.07$\pm$0.022
-Erdos 0.68$\pm$0.090(-) 0.67$\pm$0.067(-) 0.70$\pm$0.080(-) 0.50$\pm$0.028(+) 0.49$\pm$0.017 0.58$\pm$0.134(-) 0.58$\pm$0.103(-) 0.61$\pm$0.125(-) 0.28$\pm$0.026(+) 0.27$\pm$0.016
-USAir 0.73$\pm$0.114(-) 0.65$\pm$0.126(-) 0.73$\pm$0.111(-) 0.48$\pm$0.057(-) 0.41$\pm$0.045 0.70$\pm$0.168(-) 0.64$\pm$0.163(-) 0.74$\pm$0.139(-) 0.34$\pm$0.055(-) 0.27$\pm$0.050
-Netscience 0.95$\pm$0.029(-) 0.93$\pm$0.032(-) 0.96$\pm$0.026(-) 0.83$\pm$0.015(-) 0.80$\pm$0.020 0.90$\pm$0.071(-) 0.85$\pm$0.085(-) 0.91$\pm$0.068(-) 0.57$\pm$0.029(-) 0.53$\pm$0.036
-BioCelegans 0.52$\pm$0.058(-) 0.49$\pm$0.042(-) 0.49$\pm$0.067(-) 0.40$\pm$0.036($\approx$) 0.39$\pm$0.029 0.42$\pm$0.084(-) 0.37$\pm$0.073(-) 0.39$\pm$0.010(-) 0.22$\pm$0.033($\approx$) 0.21$\pm$0.024
+RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 (a) FastGreedy
   Network NMI ARI
-RAM DICE Mod SDEO MDEO RAM DICE Mod SDEO MDEO
-Dolphins 0.67$\pm$0.063(-) 0.64$\pm$0.077(-) 0.67$\pm$0.055(-) 0.57$\pm$0.031(-) 0.54$\pm$0.037 0.51$\pm$0.092(-) 0.48$\pm$0.088(-) 0.51$\pm$0.077(-) 0.30$\pm$0.029(-) 0.27$\pm$0.034
-Lesmis 0.90$\pm$0.039(-) 0.86$\pm$0.047(-) 0.90$\pm$0.044(-) 0.81$\pm$0.022(-) 0.74$\pm$0.024 0.80$\pm$0.101(-) 0.72$\pm$0.109(-) 0.77$\pm$0.120(-) 0.55$\pm$0.027(-) 0.48$\pm$0.023
-Polbooks 0.92$\pm$0.060(-) 0.88$\pm$0.044(-) 0.90$\pm$0.072(-) 0.64$\pm$0.017($\approx$) 0.63$\pm$0.014 0.93$\pm$0.064(-) 0.88$\pm$0.056(-) 0.89$\pm$0.122(-) 0.49$\pm$0.021($\approx$) 0.48$\pm$0.017
-Adjnoun 0.79$\pm$0.035(-) 0.76$\pm$0.051(-) 0.80$\pm$0.042(-) 0.72$\pm$0.022(+) 0.72$\pm$0.018 0.52$\pm$0.111(-) 0.48$\pm$0.150(-) 0.52$\pm$0.126(-) 0.20$\pm$0.035(+) 0.20$\pm$0.022
-Erdos 0.90$\pm$0.014(-) 0.87$\pm$0.022(-) 0.88$\pm$0.021(-) 0.85$\pm$0.012(+) 0.85$\pm$0.010 0.70$\pm$0.055(-) 0.62$\pm$0.071(-) 0.63$\pm$0.074(-) 0.47$\pm$0.036($\approx$) 0.45$\pm$0.031
-USAir 0.83$\pm$0.045(-) 0.82$\pm$0.066(-) 0.83$\pm$0.041(-) 0.77$\pm$0.014(-) 0.75$\pm$0.009 0.77$\pm$0.070(-) 0.74$\pm$0.139(-) 0.76$\pm$0.103(-) 0.48$\pm$0.044(-) 0.44$\pm$0.036
-Netscience 0.94$\pm$0.027(-) 0.92$\pm$0.023(-) 0.93$\pm$0.024(-) 0.86$\pm$0.010(-) 0.85$\pm$0.006 0.82$\pm$0.108(-) 0.75$\pm$0.102(-) 0.80$\pm$0.095(-) 0.53$\pm$0.026(-) 0.51$\pm$0.017
-BioCelegans 0.78$\pm$0.029(-) 0.72$\pm$0.028(-) 0.79$\pm$0.030(-) 0.74$\pm$0.011(-) 0.73$\pm$0.009 0.45$\pm$0.060(-) 0.50$\pm$0.063(-) 0.47$\pm$0.081(-) 0.31$\pm$0.017(-) 0.30$\pm$0.017
+RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO RAM DICE Mod CD-ATTACK CoeCo SDEO MDEO
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 (b) WalkTrap
@@
 * Refer to caption
-Figure 2: The evolution process of MDEO and SDEO. The attacker is FastGreedy.
-Figure 3: The evolution process of MDEO and SDEO. The attacker is WalkTrap.
-V-C Parameter sensitivity analysis
-To bolster the completeness of our study, we conduct a targeted sensitivity analysis.
-In order to validate whether the transfer contributes to the optimization, we investigate the influence of the number of networks assisting in the target network on the results, as shown in Table II.
-For fairness, the number of solutions to transfer is the same for each setting.
-As seen, the results are related to the number of assisted networks, as reflected by that the fitness values of $|\mathcal{G}^{\mathcal{S}}|=2$ and $|\mathcal{G}^{\mathcal{S}}|=3$ are larger than that of $|\mathcal{G}^{\mathcal{S}}|=1$.
-Furthermore, in the test of a total of eight networks, $|\mathcal{G}^{\mathcal{S}}|=3$ achieves the optimum on five networks, five networks, and seven networks respectively when defending against FastGreedy and WalkTrap, meaning our transfer strategy is effective in the optimization.
-We also examine the sensitivity of the ratio of elite solutions that will be directly transferred into the next generation in evolution.
-We refer this ratio as $\mu$ and set it to $\mu_{1}=0.05$, $\mu_{2}=0.10$, $\mu_{3}=0.15$ and $\mu_{4}=0.20$, respectively.
-We can observe that no ratio exhibits obvious advantages over other settings.
-For example, the large ratio of elite $\mu_{4}$ (purple line) achieves the best performance on Adjnoun, Dolphins, Lesmis, Erods and Polbooks while achieving the worst performance on USAir and BioCelegans.
-On the contrary, the setting of low ratio $\mu_{1}$, which has the best performance on USAir and BioCelegans, is inferior to other settings on the rest of networks.
-These two examples demonstrate that MDEO is not sensitive to the ratio of elite solutions.
-TABLE II: The fitness comparison of the number of networks participating in transfer in the process of optimization.
-Network FastGreedy WalkTrap
-$|\mathcal{G}^{\mathcal{S}}|=1$ $|\mathcal{G}^{\mathcal{S}}|=2$ $|\mathcal{G}^{\mathcal{S}}|=3$ $|\mathcal{G}^{\mathcal{S}}|=1$ $|\mathcal{G}^{\mathcal{S}}|=2$ $|\mathcal{G}^{\mathcal{S}}|=3$
-Dolphins 0.177 0.198 0.199 0.213 0.231 0.237
-Lesmis 0.190 0.193 0.195 0.110 0.105 0.113
-Polbooks 0.044 0.051 0.049 0.050 0.051 0.047
-Adjnoun 0.684 0.653 0.649 0.302 0.297 0.312
-Erdos 0.141 0.156 0.159 0.046 0.060 0.046
-USAir 0.031 0.038 0.040 0.023 0.022 0.024
-Netscience 0.031 0.038 0.039 0.041 0.040 0.052
-BioCelegans 0.101 0.110 0.106 0.086 0.094 0.100
-Figure 4: The fitness profile comparison of different elite transfer ratios.
-V-D Structural change
-The structural change in some indices after perturbation is shown in Table III, through which we can observe which kind of edges that the community deception algorithm prefers to modify.
+Figure 2: The optimization curve of MDEO and SDEO on eight tested networks for the illustrative task of community deception, with FastGreedy as the attacker.
+Figure 3: The optimization curve of MDEO and SDEO on eight tested networks for the illustrative task of community deception, with WalkTrap as the attacker.
+V-D Parameter sensitivity analysis
+In this section, we investigate the parameter sensitivity of the MDEO by varying three key parameters in Table IV: The transfer interval (k), the number of transfer solutions ($|\text{T}|$), and the number of assisted networks ($|\mathcal{G}^{S}|$), with the default configuration set to $\text{k}=5$, $|\text{T}|=30$, and $|\mathcal{G}^{S}|=3$.
+Compared to the default setting, increasing the transfer interval (e.g., $\text{k}=10$ and $\text{k}=20$) generally results in slightly degraded performance across most networks, indicating that more frequent knowledge transfer (lower k) is beneficial for optimization.
+Likewise, reducing the number of transfer solutions (e.g., $|\text{T}|=10$ and $|\text{T}|=20$) leads to small drops in performance when compared with the default $|\text{T}|=30$, suggesting that transferring more solutions helps guide the search more effectively.
+Furthermore, reducing the number of assisted networks (i.e., $|\mathcal{G}^{S}|=1$ or $2$) also yields worse results relative to the default, reinforcing the idea that leveraging more source networks provides richer and more useful guidance.
+We also investigate the role of community partitioning in network alignment by comparing community-based alignment with direct alignment of high- and low-degree nodes.
+The results show that community-based alignment more effectively facilitates knowledge transfer.
+The results clearly support that more frequent and comprehensive knowledge transfer leads to better optimization outcomes, indicating the effectiveness of the proposed MDEO.
+TABLE IV: Parameter sensitivity analysis of MDEO with the knowledge transfer interval k set to 5, the number of transfer solutions $|\text{T}|$ set to 30, and the number of assisted networks $|\mathcal{G}^{\mathcal{S}}|$ set to 3.
+MDEO incorporates community-based alignment by default, referred to CA.
+The tested attacker is the FastGreedy algorithm.
+The shadow cell is a better setting than default.
+  Network W/O CA $\mathbf{k=10}$ $\mathbf{k=20}$ $\mathbf{|T|=10}$ $\mathbf{|T|=20}$ $\boldsymbol{|\mathcal{G}^{\mathcal{S}}|=1}$ $\boldsymbol{|\mathcal{G}^{\mathcal{S}}|=2}$ Default
+Dolphins 0.189 0.187 0.182 0.188 0.189 0.167 0.188 0.198
+Lesmis 0.213 0.195 0.186 0.182 0.197 0.205 0.207 0.208
+Polbooks 0.044 0.048 0.041 0.038 0.047 0.044 0.051 0.051
+Adjnoun 0.644 0.623 0.621 0.578 0.633 0.644 0.653 0.649
+Erdos 0.151 0.149 0.150 0.146 0.149 0.141 0.148 0.152
+USAir 0.029 0.032 0.028 0.028 0.037 0.030 0.037 0.039
+Netscience 0.028 0.030 0.027 0.030 0.032 0.029 0.031 0.032
+BioCelegans 0.104 0.106 0.104 0.105 0.106 0.101 0.105 0.108
+Comparison 1/7 0/8 0/8 0/8 0/8 0/8 1/7 -
+As for the efficiency, Figure 4 illustrates the average running time of the MDEO for different parameter configurations related to knowledge transfer.
+The left subplot examines the impact of varying the transfer interval (k) while the right subplot evaluates different transfer solution numbers ($|\text{T}|$).
+The results indicate that both parameters have a marginal influence on the average running time.
+Specifically, as the transfer interval increases, a slight decrease in running time is observed, suggesting reduced overhead from less frequent transfers.
+Conversely, increasing the number of transfer solutions slightly increases the running time, likely due to the additional computational cost of processing more solutions.
+As observed, the impact of knowledge transfer settings on running time is minimal, indicating that the MDEO achieves an excellent balance between effectiveness and efficiency.
+Figure 4: Average running time of MDEO under different parameter settings when optimizing eight real-world networks simultaneously.
+V-E Structural change
+The structural change in some indices after perturbation is shown in Table V, through which we can observe which kind of edges that the community deception algorithm prefers to modify.
 CC refers to the clustering coefficient, measuring the extent to which nodes tend to cluster together.
@@
 The observed reduction in modularity indicates that our algorithm effectively weakens the internal connections of individual communities while simultaneously enhancing their external connectivity.
-TABLE III: The structural change after modifying edges obtained by MDEO-II. The tested attacker is FastGreedy.
-Data $|E^{+}\cup E^{-}|$ CC ASD $20\%-$ Betweenness $20\%-$ PageRank Modularity
+TABLE V: The structural modifications resulting from edge alterations performed by MDEO, with FastGreedy as the testing attacker.
+  Data $|E^{+}\cup E^{-}|$ $C$ ASD $20\%$- Betweenness $20\%$- PageRank Modularity
 Dolphins 10 $0.309\rightarrow 0.291$ $3.357\rightarrow 3.051$ $1\rightarrow 0.667$ $1\rightarrow 0.933$ $0.495\rightarrow 0.442$
@@
 USAir 100 $0.396\rightarrow 0.386$ $2.738\rightarrow 2.715$ $1\rightarrow 0.902$ $1\rightarrow 0.973$ $0.319\rightarrow 0.303$
-VI Conclusion
-In this work, we have explored a framework–multi-domain evolutionary optimization.
+V-F Generalizability analysis
+The MDEO framework, originally demonstrated through the edge-level task of community deception, is fundamentally designed for generalizability.
+It can be effectively applied to a wide range of graph-structured combinatorial problems, encompassing both node-level and edge-level objectives.
+By appropriately adapting the solution representation and fitness function to suit the specific optimization goal, MDEO is capable of addressing diverse challenges across various domains.
+As a way of illustration, we test the classical node-level task, i.e., the influence maximization task [67], where the solution sizes are set to 10 for the Email and Polblogs networks, and 15 for the Facebook and Wiki networks.
+As shown in Figures 5, MDEO consistently achieves higher average fitness values compared to SDEO, highlighting its effectiveness across different network structures.
+These representative tasks at each level collectively demonstrate the method’s generality and adaptability.
+Figure 5: The optimization curve of MDEO and SDEO in addressing the node-level problem of influence maximization.
+VI Discussion
+Although our MDEO framework has demonstrated promising performance, there still exists some challenges.
+Many real-world networks vary considerably in scale, and direct knowledge transfer across such imbalanced scenarios may introduce negative transfer effects.
+To address this issue, we observe that certain fundamental structural properties, such as modularity, centrality, and connectivity patterns, tend to remain consistent across different scales, providing a reliable foundation for effective knowledge transfer between networks of varying sizes.
+Building on this insight, we will explore extending MDEO to support a bidirectional transfer mechanism that enables adaptive knowledge exchange between large and small networks.
+This enhancement is designed to improve the robustness and adaptability of our approach in imbalanced-size scenarios, thereby enhancing its effectiveness across heterogeneous domains.
+Developing this bidirectional optimization framework is a key direction for future work and is expected to expand the practical applicability of MDEO.
+VII Conclusion
+In this work, we have explored a framework–multi-domain evolutionary optimization (MDEO).
 We actualized this concept within the realm of network structures, predicated on the inherent shared properties of real-world networks.
-The proposed MDEO consists of three components: 1) measuring the similarity of domains at the community level; 2) selecting the anchor pair nodes and training the mapping of solutions of different domains 3) optimizing different domains simultaneously and making them promote each other.
-The proposed method was validated on the task of community deception with eight different real-world networks of varying sizes from various domains, and the experimental results (with statistical analysis and sensitivity analysis) show the fitness values of evolutionary optimization on different networks have been improved, suggesting the effectiveness of our proposed multi-domain evolutionary optimization.
+The proposed method was validated on the task of community deception with eight different real-world networks of varying sizes from various domains, and the experimental results show that the fitness values of evolutionary optimization on different networks have been improved, suggesting the effectiveness of our proposed MDEO.
 As an exploratory work, MDEO has been successfully applied to networks of similar sizes.